### v5

# Binomial Transforms and the Binomial Convolution of Sequences Kunle Adegoke

Abstract
Given any two sequences of complex numbers, we establish simple relations between their binomial convolution and the binomial convolution of their individual binomial transforms. We employ these relations to derive new identities involving Fibonacci numbers, Bernoulli numbers, Catalan numbers, harmonic numbers, odd harmonic numbers, Stirling numbers of the second kind, and binomial coefficients. In addition, we present several results which allow the construction of new binomial-transform pairs from existing ones. Many new relations concerning self-inverse sequences are also derived.
2020 Mathematics Subject Classification: Primary 05A10; Secondary 05A19, 11B68, 11B39.
Keywords: Binomial transform, binomial coefficient, binomial convolution, combinatorial identity, sequence, Bernoulli number, Catalan number, harmonic number, Stirling numbers of the second kind.
1 Introduction
Let $n$ be a non-negative integer. Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be sequences of complex numbers. It is known that [8, p.4] (see also [16, p.136] and [19]):
$s_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}}\iff\sigma_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}}.$ (1.1)
Such sequences $(s_{k})$ and $(\sigma_{k})$, as given in (1.1), will be called a binomial-transform pair of the first kind.
Consider two binomial-transform pairs of the first kind, say $\{(s_{k}),(\sigma_{k})\}$ and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$. We will establish the following binomial convolution identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}s_{n-k}t_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}\tau_{n-k}}.$
Let $(\bar{s}_{k})$ and $(\bar{\sigma}_{k})$, $k=0,1,2,\ldots$, be sequences of complex numbers. It is also known that
$\bar{s}_{n}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\bar{\sigma}_{k}}\iff\bar{\sigma}_{n}=\sum_{k=0}^{n}{\binom{{n}}{k}\bar{s}_{k}}.$ (1.2)
Such sequences $(\bar{s}_{k})$ and $(\bar{\sigma}_{k})$, as given in (1.2), will be called a binomial-transform pair of the second kind. In this case we also call $(\bar{\sigma}_{k})$ the binomial transform of $(\bar{s}_{k})$.
Given two binomial-transform pairs of the second kind, say $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ and $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, we will derive the following binomial convolution identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{s}_{k}\bar{t}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{\sigma}_{k}\bar{\tau}_{n-k}}.$
Throughout this paper, we will distinguish binomial-transform pairs of the second kind with a bar on the letter representing each sequence. We will use a Greek alphabet to denote each binomial transform and the corresponding Latin alphabet to represent the original sequence.
We will also derive the following binomial convolution of one sequence and the binomial transform of the other:
$\sum_{k=0}^{n}{\binom{{n}}{k}\bar{s}_{k}\bar{\tau}_{n-k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\bar{\sigma}_{k}\bar{t}_{n-k}}.$
Given a binomial-transform pair of the first kind, $\{(s_{k}),(\sigma_{k})\}$ and a binomial-transform pair of the second kind, $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, we will derive the following binomial convolution identity:
$\sum_{k=0}^{n}{\binom{{n}}{k}s_{k}\bar{t}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}\bar{\tau}_{n-k}}.$
We will give a short proof of the known result [9, 13]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k+m}}=\sum_{k=0}^{m}{(-1)^{k}\binom{{m}}{k}\sigma_{k+n}},$
and, based partly on it, establish the following identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{n-k+m}\sum_{q=0}^{r}{(-1)^{q}\binom{{r}}{q}\tau_{k+q}}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}t_{n-k+r}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}}}.$
In fact, we will prove that
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{k+p+m}}}=\sum_{k=0}^{m}{(-1)^{k}\binom{m}{k}\sigma_{n+k+r}},$
and hence establish
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{n-k+p+m}}\sum_{j=0}^{u}{(-1)^{j}\binom{{u}}{j}t_{k+j+v}}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p+r}}\sum_{j=0}^{v}{(-1)^{j}\binom{{v}}{j}\tau_{n-k+j+u}}},$
valid for non-negative integers $m$, $n$, $r$, $u$ and $v$ and any binomial-transform pairs $\{(s_{k}),(\sigma_{k}\}$ and $\{(t_{k}),(\tau_{k}\}$, $k=0,1,2,\ldots,$ of the first kind.
Binomial transform of products of sequences will be computed in Section 5. The results will complement those of Boyadzhiev [7].
In Section 7, we will derive several results which allow one to construct new binomial-transform pairs from existing ones. For example we will show that if $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=\frac{1}{{\left({k+1}\right)\left({k+2}\right)}}\sum\limits_{j=0}^{k}{t_{j}}\text{ and }\alpha_{k}=\frac{1}{{\left({k+1}\right)\left({k+2}\right)}}\sum\limits_{j=0}^{k}{\tau_{j}}.$
If, in (1.1), $\sigma_{k}=s_{k}$ for every $k=0,1,2,\ldots$, then the sequence $(s_{k})$ is called an invariant sequence or a self-inverse sequence. If $\sigma_{k}=-s_{k}$ for every $k=0,1,2,\ldots$, then the sequence $(s_{k})$ will be called an anti-self-inverse sequence. We will show in Section 8 that every invertible sequence has a self-inverse sequence and an anti-self-inverse sequence associated with it. In particular, we will show that if the sequence $(s_{k})$, $k=0,1,2,\ldots$, is a self-inverse sequence, then the sequence $(w_{k})$ given by
$w_{k}=\frac{1}{k+1}\sum_{j=0}^{k-1}s_{j},\quad k=0,1,2,\ldots,$
is an anti-self-inverse sequence, and vice versa.
Remark 1.1.
Although conversion from one kind of binomial transform to the other is, in general, very straightforward (see Remark 12.1), it is convenient to retain the distinction between the two kinds because binomial coefficient identities occur naturally in either of the two kinds. Thus, binomial convolution identities can be written directly without the need to convert from one kind to the other.
The binomial convolution identities developed in this paper will facilitate the discovery of an avalanche of new combinatorial identities. We now present a couple of identities, selected from our results, to whet the reader’s appetite for reading on.
We derived various general identities involving binomial transform pairs, such as
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{y-k}}{x}t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{y-k}}{{y-x}}\tau_{n-k}},$
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\frac{{m\,H_{k+m}}}{{k+m}}\,t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\left({H_{k+m}-H_{k}}\right)\tau_{n-k}},$
and
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k+m}\bar{t}_{n-k}}=H_{m}\bar{\tau}_{n}-\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\frac{1}{k}\,\bar{\tau}_{n-k}},$
where $H_{k}$ is a harmonic number and $m$ is a complex number that is not a negative integer.
We discovered the following binomial convolution of harmonic and odd harmonic numbers:
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}H_{k}O_{n-k}}=\sum_{k=1}^{n-1}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2\left({n-k}\right)-1}}}{{k\left({n-k}\right)}}\binom{{2\left({n-k}\right)}}{{n-k}}^{-1}}.$
We obtained some identities involving Bernoulli numbers and Bernoulli polynomials, including the following:
[…]
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({\frac{y}{w}}\right)^{k}B_{n-k}(y)B_{k}(w)}=\frac{{ny}}{{2w}}\left({1-\left({\frac{y}{w}}\right)^{n-2}}\right)B_{n-1},\quad\text{$n$ an odd integer},$
and
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{x}}{k}B_{n-k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x+k}}{k}B_{n-k}},$
where $x$, $y$ and $w$ are complex numbers.
We found the following binomial convolution of Fibonacci and Bernoulli numbers, valid for $n$ an even integer:
$\sum_{k=0}^{n}{\binom{{n}}{k}F_{k}B_{n-k}}=0.$
We obtained the following polynomial identities involving, respectively, harmonic numbers and odd harmonic numbers:
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k}t^{k}}=\sum_{k=1}^{n}{(-1)^{k-1}\binom{{n}}{k}\frac{1}{k}\,t^{k}\left({1+t}\right)^{n-k}}$
and
$\sum_{k=0}^{n}{\binom{{n}}{k}O_{k}t^{k}}=\sum_{k=1}^{n}{(-1)^{k-1}\binom{{n}}{k}2^{2k-1}\binom{2k}{k}^{-1}\frac{1}{k}\,t^{k}\left({1+t}\right)^{n-k}},$
with the special values
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k}}=\sum_{k=1}^{n}{(-1)^{k-1}2^{n-k}\binom{{n}}{k}\frac{1}{k}}$
and
$\sum_{k=0}^{n}{\binom{{n}}{k}O_{k}}=\sum_{k=1}^{n}{(-1)^{k-1}2^{n+k-1}\binom{{n}}{k}\binom{2k}{k}^{-1}\frac{1}{k}}.$
Let
$C_{n}=\frac{1}{n+1}\binom{2n}{n},\quad n=0,1,2,\ldots,$
be the Catalan numbers. By employing a result of Mikić [18], we derived the following identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{{\left\lfloor{k/2}\right\rfloor}}\binom{{n-k}}{{\left\lfloor{\left({n-k}\right)/2}\right\rfloor}}}=\begin{cases}C_{n/2}\binom{n}{n/2},&\text{$n$ even};\\ 0,&\text{$n$ odd};\end{cases}$
where here and throughout this paper, $\lfloor x\rfloor$ is the greatest integer less than or equal to $x$.
We derived a couple of binomial coefficient identities, including the following:
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{n-k+m}}{u}\binom{{r}}{{j-k}}}=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{n-k+r}}{j}\binom{{m}}{{u-k}}}$,
which holds for complex numbers $j$, $m$, $r$, and $u$.
We derived many, mostly new, binomial transform identities involving special numbers, to serve as a pool of binomial-transform pairs. These are presented in Section 13.
We close this section by giving a short description of each of the special numbers from which illustrations will be drawn.
Binomial coefficients are defined, for non-negative integers $i$ and $j$, by
$\binom{i}{j}=\begin{cases}\dfrac{{i!}}{{j!(i-j)!}},&\text{$i\geq j$};\\ 0,&\text{$i<j$};\end{cases}$
the number of distinct sets of $j$ objects that can be chosen from $i$ distinct objects.
Generalized binomial coefficients are defined for complex numbers $r$ and $s$ by
$\binom{r}{s}=\frac{{\Gamma(r+1)}}{{\Gamma(s+1)\Gamma(r-s+1)}},$ (1.3)
where the Gamma function, $\Gamma(z)$, is defined for $\Re(z)>0$ by
$\Gamma(z)=\int_{0}^{\infty}{e^{-t}t^{z-1}dt}=\int_{0}^{\infty}{\left({\log(1/t)}\right)^{z-1}dt},$
and is extended to the rest of the complex plane, excluding the non-positive integers, by analytic continuation.
Harmonic numbers, $H_{n}$, and odd harmonic numbers, $O_{n}$, are defined for non-negative integers $n$ by
$H_{n}=\sum_{k=1}^{n}\frac{1}{k},\quad O_{n}=\sum_{k=1}^{n}\frac{1}{2k-1},\quad H_{0}=0=O_{0}.$
The Bernoulli numbers, $B_{k}$, are defined by the generating function
$\frac{z}{{e^{z}-1}}=\sum_{k=0}^{\infty}{B_{k}\frac{{z^{k}}}{{k!}}},\quad z<2\pi\,$,
and the Bernoulli polynomials by the generating function
$\frac{{ze^{xz}}}{{e^{z}-1}}=\sum_{k=0}^{\infty}{B_{k}(x)\frac{{z^{k}}}{{k!}}},\quad|z|<2\pi\,.$
Clearly, $B_{k}=B_{k}(0)$.
The first few Bernoulli numbers are
$B_{0}=1,\,B_{1}=-\frac{1}{2},\,B_{2}=\frac{1}{6},\,B_{3}=0,\,B_{4}=-\frac{1}{{30}},\,B_{5}=0,\,B_{6}=\frac{1}{{42}},\,B_{7}=0,\,\ldots\,$,
while the first few Bernoulli polynomials are
$\begin{split}&B_{0}(x)=1,\quad B_{1}(x)=x-\frac{1}{2},\quad B_{2}(x)=x^{2}-x+\frac{1}{6},\quad B_{3}(x)=x^{3}-\frac{3}{2}x^{2}+\frac{1}{2}x.\end{split}$
An explicit formula for the Bernoulli polynomials is
$\frac{B_{n}(x)}{x^{n}}=\sum_{k=0}^{n}{\binom{n}{k}\frac{B_{k}}{x^{k}}}\,$,
while a recurrence formula for them is
$B_{n}(x+1)=\sum_{k=0}^{n}{\binom{n}{k}B_{k}(x)}.$ (1.4)
The Fibonacci numbers, $F_{n}$, and the Lucas numbers, $L_{n}$, are defined, for $n\in\mathbb{Z}$, through the recurrence relations
$F_{n}=F_{n-1}+F_{n-2},\mbox{($n\geq 2$)},\quad\mbox{$F_{0}=0$, $F_{1}=1$};$
and
$L_{n}=L_{n-1}+L_{n-2},\mbox{($n\geq 2$)},\quad\mbox{$L_{0}=2$, $L_{1}=1$};$
with
$F_{-n}=(-1)^{n-1}F_{n}\,,\quad L_{-n}=(-1)^{n}L_{n}\,.$ (1.5)
Explicit formulas (Binet formulas) for the Fibonacci and Lucas numbers are
$F_{n}=\frac{{\alpha^{n}-\beta^{n}}}{{\alpha-\beta}},\quad L_{n}=\alpha^{n}+\beta^{n},\quad n\in\mathbb{Z},$
where $\alpha=(1+\sqrt{5})/2$ is the golden ratio and $\beta=(1-\sqrt{5})/2=-1/\alpha$.
The generalized Fibonacci sequence, $(G_{n})$, $n=0,1,2,\ldots$, the so-called gibonacci sequence, is a generalization of $F_{n}$ and $L_{n}$; having the same recurrence relation as the Fibonacci sequence but with arbitrary initial values. Thus
$G_{n}=G_{n-1}+G_{n-2},\quad(n\geq 2);$
with
$\quad G_{-n}=G_{-n+2}-G_{-n+1}.$
where $G_{0}$ and $G_{1}$ arbitrary numbers (usually integers) not both zero.
Stirling numbers of the second kind are defined for non-negative integers $m$ and $n$ through the generating function
[…]
and hence by binomial inversion, they have the explicit representation
[…]
The Stirling numbers of the second kind obey the following recursion relation [15, Equation (6.3), p.259]:
[…]
and have the following special values:
[…]
and
[…]
2 Identities involving binomial-transform pairs of the first kind
Our first main result is stated in Theorem 2.1; we require the result stated in the next lemma for its proof.
Lemma 2.1.
Let $\left\{(a_{k}),(\alpha_{k})\right\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. Let $\mathcal{L}_{x}$ be a linear operator defined by $\mathcal{L}_{x}(x^{j})=a_{j}$ for every complex number $x$ and every non-negative integer $j$. Then $\mathcal{L}_{x}((1-x)^{j})=\alpha_{j}$.
Proof.
We have
$\mathcal{L}_{x}\left({\left({1-x}\right)^{j}}\right)=\mathcal{L}_{x}\left({\sum_{k=0}^{j}{(-1)^{k}\binom{{j}}{k}x^{k}}}\right)$ $=\sum_{k=0}^{j}{(-1)^{k}\binom{{j}}{k}\mathcal{L}_{x}\left({x^{k}}\right)}$
$=\sum_{k=0}^{j}{(-1)^{k}\binom{{j}}{k}a_{k}}$
$=\alpha_{j}.$
Theorem 2.1.
Let $n$ be a non-negative integer. If $\{(s_{k}),(\sigma_{k})\}$ and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the first kind, then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}s_{k}t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}\tau_{n-k}}.$ (2.1)
Proof.
Let $n$ be a non-negative integer. Let $x$ and $y$ be complex numbers. Consider the following identity whose veracity is readily established by the binomial theorem:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{n-k}x^{k}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\left({1-y}\right)^{n-k}\left({1-x}\right)^{k}}.$ (2.2)
Let $(t_{j})$, $j=0,1,2,\ldots$, be a sequence of complex numbers. Let $\mathcal{L}_{x}(x^{j})=t_{j}$ .
Operate on both sides of (2.2) with $\mathcal{L}_{x}$ to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{n-k}t_{k}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\left({1-y}\right)^{n-k}\tau_{k}},$
where
$\tau_{k}=\mathcal{L}_{x}((1-x)^{k})=\sum_{i=0}^{k}{(-1)^{i}\binom{{k}}{i}t_{i}}.$
Thus,
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}y^{k}t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({1-y}\right)^{k}\tau_{n-k}}.$ (2.3)
Let $(s_{j})$, $j=0,1,2,\ldots$, be a sequence of complex numbers. Let $\mathcal{L}_{y}(y^{j})=s_{j}$. The action of $\mathcal{L}_{y}$ on (2.3) produces
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}s_{k}t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}\tau_{n-k}},$
where
$\sigma_{k}=\mathcal{L}_{y}((1-y)^{k})=\sum_{i=0}^{k}{(-1)^{i}\binom{{k}}{i}s_{i}},$
and the proof is complete. ∎
In the next example we illustrate Theorem 2.1.
Example 2.1.
Consider the following identity [8, Equation (10.7)], where $x$ and $y$ are complex numbers and $n$ is a non-negative integer:
$\sum_{k=0}^{n}{\binom{{n}}{k}(-1)^{k}\binom{{y-k}}{x}}=\binom{{y-n}}{{y-x}}$. (2.4)
We can choose
$s_{k}=\binom{{y-k}}{x},\quad\sigma_{k}=\binom{{y-k}}{{y-x}}$,
and use these in (2.1) to obtain the following identity
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{y-k}}{x}t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{y-k}}{{y-x}}\tau_{n-k}},$ (2.5)
which holds for every binomial-transform pair of the first kind $\{(t_{k}),(\tau_{k})\}$.
As an immediate consequence of (2.5), we have the following polynomial identity in the complex variable $t$:
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{y-k}}{x}t^{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{y-k}}{{y-x}}\left({1-t}\right)^{n-k}},$
which can also be written as
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{y+k}}{x}t^{k}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{y+k}}{{y+n-x}}\left({1-t}\right)^{k}},$ (2.6)
which is valid for all complex numbers $x$, $y$ and $t$.
Example 2.2.
If we identify
$s_{k}=\frac{{H_{k+m}}}{{k+m}},\quad\sigma_{k}=\frac{{H_{k+m}-H_{k}}}{m}\binom{{k+m}}{k}^{-1},$
from the following identity [8, Equation (9.46)]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{H_{k+m}}}{{k+m}}}=\frac{{H_{n+m}-H_{n}}}{m}\binom{{n+m}}{n}^{-1},\quad m\neq 0,$
then from (2.1), we obtain the following identity
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\frac{{m\,H_{k+m}}}{{k+m}}\,t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\left({H_{k+m}-H_{k}}\right)\tau_{n-k}},$ (2.7)
which holds for any binomial-transform pair of the first kind, $\{(t_{k}),(\tau_{k})\}$, for every complex number $m$ that is not a non-positive number and every non-negative integer $n$.
In particular, since
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{G_{tk+r}}}{{L_{t}^{k}}}}=\frac{{(-1)^{r}}}{{L_{t}^{n}}}\left({G_{0}L_{tn-r}-G_{tn-r}}\right),$ (2.8)
using
$t_{k}=\frac{{G_{tk+r}}}{{L_{t}^{k}}},\quad\tau_{k}=\frac{{(-1)^{r}}}{{L_{t}^{k}}}\left(G_{0}L_{tk-r}-G_{tk-r}\right),$
in (2.7) yields the following general Fibonacci-harmonic number identity:
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\frac{{m\,H_{k+m}}}{{k+m}}L_{t}^{k}G_{t(n-k)+r}},$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\left({H_{k+m}-H_{k}}\right)L_{t}^{k}\left({G_{0}L_{t(n-k)-r}-G_{t(n-k)-r}}\right)};$ (2.9)
and, in particular,
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\frac{{m}}{{k+m}}H_{k+m}F_{n-k}}=\sum_{k=0}^{n}{(-1)^{k-1}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\left({H_{k+m}-H_{k}}\right)F_{n-k}},$ (2.10)
with the special value
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\frac{{H_{k+1}F_{n-k}}}{{k+1}}}=\sum_{k=0}^{n}{(-1)^{k-1}\binom{{n}}{k}\frac{{F_{n-k}}}{{\left({k+1}\right)^{2}}}}.$ (2.11)
Example 2.3.
From (2.4) and [8, Equation (10.1)]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{x}}{k}\binom{{y}}{k}^{-1}}=\binom{{y-x}}{n}\binom{{y}}{n}^{-1},\quad y-n\not\in\mathbb{Z}^{-},y\in\mathbb{C},$ (2.12)
we can choose
$s_{k}=\binom{{y-k}}{x},\quad\sigma_{k}=\binom{{y-k}}{{y-x}}$,
and
$t_{k}=\binom{{u}}{k}\binom{{v}}{k}^{-1},\quad\tau_{k}=\binom{{v-u}}{k}\binom{{v}}{k}^{-1}.$
Using these in (2.1), we obtain the following identity
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{y-k}}{x}\binom{u}{{n-k}}\binom{{v}}{{n-k}}^{-1}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{y-k}}{{y-x}}\binom{{v-u}}{{n-k}}\binom{{v}}{{n-k}}^{-1}},$ (2.13)
which is valid for complex numbers $x$, $y$, $u$ and $v$ such that vnv-n is not a negative integer.
In particular, at $v=n$, we get
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{y-k}}{x}\binom{u}{{n-k}}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{y-k}}{{y-x}}\binom{{n-u}}{{n-k}}}$, (2.14)
with the special value
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}\binom{{y-k}}{{y-x}}}=\binom{y-n}{x}$, (2.15)
which is the binomial transform of the first kind of (2.4).
Example 2.4.
From [8, Equation (9.37)]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}H_{k+m}}=-\frac{1}{n}\binom{{n+m}}{m}^{-1},\quad n\neq 0,$ (2.16)
and [8, Equation (9.51)]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}O_{k}}=-\binom{{2n}}{n}^{-1}\frac{{2^{2n-1}}}{n},$ (2.17)
we can identify
$s_{k}=H_{k+m},\quad\sigma_{k}=\frac{{\delta_{k0}(1+H_{m})-1}}{{k+\delta_{k0}}}\binom{{k+m}}{m}^{-1},$
and
$t_{k}=O_{k},\quad\tau_{k}=-\frac{{1-\delta_{k0}}}{{k+\delta_{k0}}}\binom{{2k}}{k}^{-1}2^{2k-1}.$
Here and throughout this paper, $\delta_{ij}$ is Kronecker’s delta having the value $1$ when $i=j$ and zero otherwise.
Using these in (2.1) leads to the following harmonic number identity:
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}H_{k+m}O_{n-k}}$
$\qquad=-\frac{H_{m}}{n}\binom{2n}{n}^{-1}2^{2n-1}+\sum_{k=1}^{n-1}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2\left({n-k}\right)-1}}}{{k\left({n-k}\right)}}\binom{{k+m}}{m}^{-1}\binom{{2\left({n-k}\right)}}{{n-k}}^{-1}},$ (2.18)
which is valid for every positive integer $n$ and every complex number $m$ that is not a negative integer.
In particular, we have the following alternating binomial convolution of harmonic number and odd harmonic number:
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}H_{k}O_{n-k}}=\sum_{k=1}^{n-1}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2\left({n-k}\right)-1}}}{{k\left({n-k}\right)}}\binom{{2\left({n-k}\right)}}{{n-k}}^{-1}}.$ (2.19)
Corollary 2.2.
Let $n$ be a non-negative integer. If $\{(s_{k}),(\sigma_{k})\}$ and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the first kind, then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\sigma_{k}t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}\tau_{n-k}}.$ (2.20)
Corollary 2.3.
Let $n$ be a non-negative integer. If $\{(s_{k}),(\sigma_{k})\}$, $k=0,1,2,\ldots$, is a binomial-transform pair of the first kind, then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}s_{k}s_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}\sigma_{n-k}}.$ (2.21)
Remark 2.1.
We deduce from (2.21) that if $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}s_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}\sigma_{n-k}},$ (2.22)
with each sum being 0 for odd $n$.
Example 2.5.
In view of the following identity [23]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{k}}{{\left\lfloor{k/2}\right\rfloor}}}=2^{-n}C_{n},$ (2.23)
we can choose
$s_{k}=2^{-k}\binom{{k}}{{\left\lfloor{k/2}\right\rfloor}},\quad\sigma_{k}=2^{-k}C_{k},$ (2.24)
and use in (2.21) to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{{\left\lfloor{k/2}\right\rfloor}}\binom{{n-k}}{{\left\lfloor{\left({n-k}\right)/2}\right\rfloor}}}=(-1)^{n}\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}C_{k}C_{n-k}}.$ (2.25)
Mikić [18] has shown that if $n$ is a non-negative integer, then
[…]
From (2.25) and (2.26), we obtain
[…]
Example 2.6.
Plugging $s_{k}$ and $\sigma_{k}$ from Example 2.2 into (2.22) yields, for $n$ a non-negative integer and $m$ a complex number that is not a negative integer,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{m^{2}H_{k+m}H_{n-k+m}}}{{\left({k+m}\right)\left({n-k+m}\right)}}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{k}^{-1}\binom{{n-k+m}}{{n-k}}^{-1}\left({H_{k+m}-H_{k}}\right)\left({H_{n-k+m}-H_{n-k}}\right)},$ (2.28)
and, in particular,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{H_{k+1}H_{n-k+1}}}{{\left({k+1}\right)\left({n-k+1}\right)}}}=\sum_{k=0}^{n}{\frac{{(-1)^{k}}}{{\left({k+1}\right)^{2}\left({n-k+1}\right)^{2}}}\binom{{n}}{k}}$. (2.29)
Example 2.7.
From (2.12), with
$s_{k}=\binom{{x}}{k}\binom{{y}}{k}^{-1},\quad\sigma_{k}=\binom{{y-x}}{k}\binom{{y}}{k}^{-1},$
identity (2.22) gives
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{x}}{k}\binom{{x}}{{n-k}}\binom{y}{k}^{-1}\binom{{y}}{{n-k}}^{-1}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{y-x}}{k}\binom{{y-x}}{{n-k}}\binom{{y}}{k}^{-1}\binom{{y}}{{n-k}}^{-1}},$ (2.30)
which, upon setting $x=n$ and $y=n$, in turn, yields
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}^{3}\binom{y}{k}^{-1}\binom{{y}}{{n-k}}^{-1}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{y-n}}{k}\binom{{y-n}}{{n-k}}\binom{{y}}{k}^{-1}\binom{{y}}{{n-k}}^{-1}}$ (2.31)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{x}}{k}\binom{{x}}{{n-k}}\binom{{n}}{k}^{-1}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n-x}}{k}\binom{{n-x}}{{n-k}}\binom{{n}}{k}^{-1}}.$ (2.32)
Setting y=1y=-1 in (2.7) gives, after some simplification,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}^{3}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{n+k}}{k}\binom{{2n-k}}{{n-k}}}$,
the left hand side of which is known, Dixon’s identity [10]:
[…]
We therefore obtain
[…]
Similarly, setting x=1x=-1 in (2.32) gives
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}^{-1}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{\left({n+1}\right)^{2}}}{{\left({n-k+1}\right)\left({k+1}\right)}}},$
from which we get
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{1}{{\left({n-k+1}\right)\left({k+1}\right)}}}=\frac{{1+(-1)^{n}}}{{\left({n+1}\right)\left({n+2}\right)}},$
upon using [28]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}^{-1}}=\left({1+(-1)^{n}}\right)\frac{{n+1}}{{n+2}},$
and hence,
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}\frac{1}{k+1}}=\frac{1}{n+1}.$
A sequence $(s_{k})$, $k=0,1,2,\ldots$, for which
$s_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}},$
for every non-negative integer $n$ will be called a self-inverse sequence, following Sun [26].
A sequence $(s_{k})$, $k=0,1,2,\ldots$, for which
$s_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}},$
for every non-negative integer $n$ will be called an anti-self-inverse sequence. Many self-inverse sequences are known (see e.g. Sun [26], Chen [9], Boyadzhiev [8]). Examples include $(L_{k})$, $(kF_{k-1})$, $(\binom{x/2}{k}/\binom{x}{k})$, $0\neq x\not\in\mathbb{Z}^{+}$, $(\binom{2k}{k}/2^{2k})$. Anti-self-inverse sequences include $(F_{k})$ and $(H_{k}/(k+1))$. Self-inverse and anti-self-inverse sequences are discussed in more detail in Section 8.
Corollary 2.4.
Let $(s_{k})$ and $(t_{k})$, $k=0,1,2,\ldots$, be two sequences of complex numbers. Let $n$ be a non-negative integer. If both sequences are invariant or both are anti-self-inverse, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}t_{n-k}}=0,\text{ if $n$ is odd;}$ (2.35)
while if one of the sequences is invariant and the other is anti-self-inverse, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}t_{n-k}}=0,\text{ if $n$ is even.}$ (2.36)
Remark $2.$
The results stated in Corollary 2.4 were also obtained by Wang [31].
Example 2.8.
If $n$ is a non-negative odd integer and $x$ is a non-zero complex number such that xnx-n is not a negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{x/2}}{k}\binom{{x}}{k}^{-1}L_{n-k}}=0.$ (2.37)
3 Identities involving binomial-transform pairs of the second kind
Lemma 3.1.
Let $\left\{(\bar{a}_{k}),(\bar{\alpha}_{k})\right\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the second kind. Let $\mathcal{M}_{x}$ be a linear operator defined by $\mathcal{M}_{x}(x^{j})=\bar{a}_{j}$ for every complex number $x$ and every non-negative integer $j$. Then $\mathcal{M}_{x}((1+x)^{j})=\bar{\alpha}_{j}$.
Proof.
We have
$\mathcal{M}_{x}\left({\left({1+x}\right)^{j}}\right)=\mathcal{M}_{x}\left({\sum_{k=0}^{j}{\binom{{j}}{k}x^{k}}}\right)$ $=\sum_{k=0}^{j}{\binom{{j}}{k}\mathcal{M}_{x}\left({x^{k}}\right)}$
$=\sum_{k=0}^{j}{\binom{{j}}{k}\bar{a}_{k}}$
$=\bar{\alpha}_{j}.$
Theorem 3.1.
Let $n$ be a non-negative integer. If $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ and $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the second kind, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{s}_{k}\bar{t}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{\sigma}_{k}\bar{\tau}_{n-k}}.$ (3.1)
Proof.
Write (2.2) as
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{n-k}x^{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({1+y}\right)^{n-k}\left({1+x}\right)^{k}}.$ (3.2)
Let $(\bar{t}_{j})$, $j=0,1,2,\ldots$, be a sequence of complex numbers. Let $\mathcal{M}_{x}(x^{j})=\bar{t}_{j}$ .
Operate on both sides of (3.2) with $\mathcal{M}_{x}$ to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{n-k}\bar{t}_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({1+y}\right)^{n-k}\bar{\tau}_{k}},$
where
$\bar{\tau}_{k}=\mathcal{M}_{x}((1+x)^{k})=\sum_{i=0}^{k}{\binom{{k}}{i}\bar{t}_{i}}.$
Thus,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{k}\bar{t}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({1+y}\right)^{k}\bar{\tau}_{n-k}}.$ (3.3)
Let $(\bar{s}_{j})$, $j=0,1,2,\ldots$, be a sequence of complex numbers. Let $\mathcal{M}_{y}(y^{j})=\bar{s}_{j}$. The action of $\mathcal{M}_{y}$ on (3.3) produces
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{s}_{k}\bar{t}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{\sigma}_{k}\bar{\tau}_{n-k}},$
where
$\bar{\sigma}_{k}=\mathcal{M}_{y}((1+y)^{k})=\sum_{i=0}^{k}{\binom{{k}}{i}\bar{s}_{i}},$
and the proof is complete. ∎
Example 3.1.
Consider the general recurrence relation for Bernoulli polynomials:
$\frac{{B_{n}\left({x+y}\right)}}{{y^{n}}}=\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k}(x)}}{{y^{k}}}}.$ (3.4)
If we choose
$\bar{s}_{k}=\frac{{B_{k}(x)}}{{y^{k}}},\quad\bar{\sigma}_{k}=\frac{{B_{k}\left({x+y}\right)}}{{y^{k}}},$
in (3.1), we obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{k}B_{n-k}(x)\bar{t}_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{k}B_{n-k}\left({x+y}\right)\bar{\tau}_{k}},$ (3.5)
which is valid for complex variables $x$ and $y$ and any binomial transform pair $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ of the second kind.
Many results can be derived from (3.5). To begin with, we have the following polynomial identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{k}B_{n-k}(x)t^{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}y^{k}B_{n-k}\left({x+y}\right)\left({1+t}\right)^{k}},$ (3.6)
which holds for complex numbers $x$, $y$ and $t$.
Choosing
$\bar{t}_{k}=\frac{{B_{k}(z)}}{{w^{k}}},\quad\bar{\tau}_{k}=\frac{{B_{k}\left({z+w}\right)}}{{w^{k}}},$
in (3.5) leads to
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({y/w}\right)^{k}B_{n-k}(x)B_{k}(z)}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({y/w}\right)^{k}B_{n-k}\left({x+y}\right)B_{k}\left({z+w}\right)},$ (3.7)
giving, in particular,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}B_{n-k}(x)B_{k}(z)}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}B_{n-k}\left({x+w}\right)B_{k}\left({z+w}\right)},$ (3.8)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({y/w}\right)^{k}B_{n-k}\left({y}\right)B_{k}\left({w}\right)}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({y/w}\right)^{k}B_{n-k}B_{k}}.$ (3.9)
Proposition 3.1.
If $n$ is a positive odd integer and $y$ and $w$ are complex numbers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({\frac{y}{w}}\right)^{k}B_{n-k}(y)B_{k}(w)}=\frac{{ny}}{{2w}}\left({1-\left({\frac{y}{w}}\right)^{n-2}}\right)B_{n-1}.$ (3.10)
Proof.
This is a consequence of the fact that the right-hand side of (3.9) can be written as
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}(y/w)^{2k}B_{n-2k}B_{2k}}-\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{\binom{{n}}{{2k-1}}(y/w)^{2k-1}B_{n-2k+1}B_{2k-1}},$
and the fact that $B_{j}$ is zero for odd $j$ greater than unity. ∎
Example 3.2.
Using the binomial-transform pairs
$\bar{s}_{k}=\binom{{x}}{k},\quad\bar{\sigma}_{k}=\binom{{x+k}}{k}$, (3.11)
obtained from [8, Equation (10.10d)]:
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{k}}=\binom{{n+x}}{n}$ (3.12)
and
$\bar{t}_{k}=B_{k},\quad\bar{\tau}_{k}=(-1)^{k}B_{k},$ (3.13)
found from the recurrence relation
$\sum_{k=0}^{n}{\binom{n}{k}B_{k}}=(-1)^{n}B_{n},$
in (3.1) gives
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{x}}{k}B_{n-k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x+k}}{k}B_{n-k}},$ (3.14)
which holds for every non-negative integer $n$ and every complex number $x$.
Proposition 3.2.
If $n$ is a non-negative integer and $x$ is a complex number such that xnx-n is not a negative integer, then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{x}}{k}\left({H_{x}-H_{x-k}}\right)B_{n-k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x+k}}{k}\left({H_{x+k}-H_{x}}\right)B_{n-k}}.$ (3.15)
In particular,
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{2k}}{k}2^{-2k}O_{k}B_{n-k}}=0,\quad\text{$n$ an even integer}.$ (3.16)
Proof.
Differentiate (3.14) wth respect to $x$, using
$\frac{d}{{dx}}\binom{{x+a}}{k}=\binom{{x+a}}{k}\left({H_{x+a}-H_{x+a-k}}\right).$
In deriving (3.16), we set x=1/2x=-1/2 in (3.15) and used
$H_{k-1/2}=2O_{k}-2\ln 2,$ (3.17)
$\binom{{-1/2}}{k}=(-1)^{k}\binom{{2k}}{k}2^{-2k},$ (3.18)
and
$\binom{{k-1/2}}{k}=\binom{{2k}}{k}2^{-2k}.$ (3.19)
Corollary 3.2.
Let $n$ be a non-negative integer. If $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ is a binomial-transform pair of the second kind, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{s}_{k}\bar{s}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{\sigma}_{k}\bar{\sigma}_{n-k}}.$ (3.20)
Example 3.3.
Using, in (3.20), the $\bar{s}_{k}$ and $\bar{\sigma}_{k}$ given in (3.11), we have
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{x}}{k}\binom{{x}}{{n-k}}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{x+k}}{k}\binom{{x+n-k}}{{n-k}}}$; (3.21)
which holds for every non-negative integer $n$ and every complex number $x$.
Example 3.4.
From [8, Equation (9.62)]:
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{k}H_{k}}=\binom{{n+m}}{m}\left({H_{m}+H_{n}-H_{m+n}}\right),$ (3.22)
we identify the following binomial-transform pair of the second kind:
$\bar{s}_{k}=\binom{{m}}{k}H_{k},\quad\bar{\sigma}_{k}=\binom{{k+m}}{m}\left({H_{m}+H_{k}-H_{k+m}}\right),$ (3.23)
which when used in (3.20) gives
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{m}}{k}\binom{{m}}{{n-k}}H_{k}H_{n-k}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{m}\binom{{n-k+m}}{m}\left({H_{m}+H_{k}-H_{k+m}}\right)\left({H_{m}+H_{n-k}-H_{n-k+m}}\right)},$ (3.24)
which is valid for every non-negative integer $n$ and every complex number $m$ that is not a negative integer.
Theorem 3.3.
Let $n$ be a non-negative integer. If $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ and $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the second kind, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\bar{s}_{k}\bar{\tau}_{n-k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\bar{\sigma}_{k}\bar{t}_{n-k}}.$ (3.25)
Proof.
Consider the following variation on (3.2):
$\sum_{k=0}^{n}{\binom{{n}}{k}y^{n-k}\left({1+x}\right)^{k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\left({1+y}\right)^{n-k}x^{k}},$ (3.26)
and proceed as in the proof of Theorem 3.1. ∎
Example 3.5.
Using the binomial-transform pairs of the second kind
[…]
and
[…]
in (3.25) yields
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{k}\binom{{x+n-k}}{{n-k}}H_{k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{{n-k}}\binom{{k+m}}{m}\left({H_{m}+H_{k}-H_{k+m}}\right)}.$ (3.27)
Example 3.6.
Consider the following identity [29]:
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+k}}^{-1}}=\frac{{m+1}}{{m-n+1}}\binom{{m-n}}{p}^{-1},$ (3.28)
which holds for every non-negative integer $n$ and complex numbers $m$ and $p$ for which mnpm-n-p is not a negative integer.
Choosing
$\bar{s}_{k}=\binom{{m}}{{p+k}}^{-1},\quad\bar{\sigma}_{k}=\frac{{m+1}}{{m-k+1}}\binom{{m-k}}{p}^{-1},$ (3.29)
in (3.25) gives
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+k}}^{-1}\bar{\tau}_{n-k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{m+1}}{{m-k+1}}\binom{{m-k}}{p}^{-1}t_{n-k}},$
which can also be written as
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+n-k}}^{-1}\bar{\tau}_{k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{m+1}}{{m-n+k+1}}\binom{{m-n+k}}{p}^{-1}t_{k}}.$ (3.30)
Example 3.7.
Differentiating (3.28) with respect to $m$ using
$\frac{d}{{dm}}\binom{{m}}{{p+k}}^{-1}=\binom{{m}}{{p+k}}^{-1}\left({H_{m-p-k}-H_{m}}\right)$
gives
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+k}}^{-1}\left({H_{m-p-k}-H_{m}}\right)}$
$\qquad=\frac{{-n}}{{\left({m-n+1}\right)^{2}}}\binom{{m-n}}{p}^{-1}-\frac{{m+1}}{{m-n+1}}\binom{{m-n}}{p}^{-1}\left({H_{m-n}-H_{m-p-n}}\right),$ (3.31)
for $m$ and $p$ complex numbers such that $m-n-p\neq\mathbb{Z}^{-}$, which at $p=0$ gives
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{k}^{-1}\left({H_{m-k}-H_{m}}\right)}=\frac{{-n}}{{\left({m-n+1}\right)^{2}}},$ (3.32)
which yields
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{k}^{-1}H_{m-k}}=\frac{{m+1}}{{m-n+1}}H_{m}-\frac{n}{{\left({m-n+1}\right)^{2}}},$
of which the well-known identity
$\sum_{k=0}^{n}{H_{k}}=\left({n+1}\right)H_{n}-n,$ (3.33)
is a particular case.
Proposition 3.3.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{2k}}{k}^{-1}2^{2k}O_{k}}=\frac{{-2n}}{{\left({2n-1}\right)^{2}}}.$ (3.34)
Proof.
Set m=1/2m=-1/2 in (3.32) and use (3.17) and (3.18). ∎
Example 3.8.
Differentiating (3.28) with respect to $p$ using
$\frac{d}{{dp}}\binom{{m}}{{p+k}}^{-1}=\binom{{m}}{{p+k}}^{-1}\left({H_{p+k}-H_{m-p-k}}\right)$
yields
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+k}}^{-1}\left({H_{m-p-k}-H_{p+k}}\right)}=\frac{{m+1}}{{m-n+1}}\binom{{m-n}}{p}^{-1}\left({H_{m-p-n}-H_{p}}\right),$ (3.35)
which can also be written as
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+n-k}}^{-1}\left({H_{m-p-n+k}-H_{p+n-k}}\right)}$
$\qquad=\frac{{m+1}}{{m-n+1}}\binom{{m-n}}{p}^{-1}\left({H_{m-p-n}-H_{p}}\right).$ (3.36)
Using
$\bar{s}_{k}=\binom{{m}}{{p+n-k}}^{-1}\left({H_{m-p-n+k}-H_{p+n-k}}\right),\quad\bar{\sigma}_{k}=\frac{{m+1}}{{m-k+1}}\binom{{m-k}}{p}\left({H_{m-p-k}-H_{p}}\right),$
in (3.25) leads to the following identity:
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+n-k}}^{-1}\left({H_{m-p-n+k}-H_{p+n-k}}\right)\bar{\tau}_{k}}$
$\qquad=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m-n+k}}{p}^{-1}\frac{{m+1}}{{m-n+k+1}}\left({H_{m-p-n+k}-H_{p}}\right)\bar{t}_{k}},$ (3.37)
which holds for an arbitrary binomial-transform pair of the second kind and complex numbers $m$ and $p$ for which mnpm-n-p is not a negative integer.
Proposition 3.4.
If $n$ is a non-negative integer and $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, is a binomial-transform pair of the second kind, then
$\sum_{k=0}^{n}{\binom{{n+1}}{{k+1}}H_{k}\bar{t}_{k}}=\sum_{k=0}^{n}{\left({H_{k}-H_{n-k}}\right)\bar{\tau}_{k}}.$ (3.38)
In particular,
$\sum_{k=0}^{n}{\binom{{n+1}}{{k+1}}H_{k}B_{k}}=\sum_{k=0}^{n}{(-1)^{k}\left({H_{k}-H_{n-k}}\right)B_{k}}$ (3.39)
and
$\sum_{k=0}^{n}{\binom{{n+1}}{{k+1}}H_{k}x^{k}}=\sum_{k=0}^{n}{\left({H_{k}-H_{n-k}}\right)\left({1+x}\right)^{k}},$ (3.40)
for every complex number $x$.
Proof.
Set $m=n$ and $p=0$ in (3.8). ∎
By considering the binomial-transform pair of the second kind $\{(1),(2^{k})\}$, $k=0,1,2,\ldots$ and making use of Theorem 3.3, we obtain the next result expressing the binomial transform of a binomial transform.
Theorem 3.4.
Let $n$ be a non-negative integer. If $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ is a binomial-transform pair of the second kind, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\bar{\sigma}_{k}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{n-k}\bar{s}_{k}}.$ (3.41)
Remark 3.1.
Identity (3.41) corresponds to Example 4 of Boyadzhiev [7].
Example 3.9.
With [8, Equation (10.10a)]
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{{k+z}}}=\binom{{n+x}}{{n+z}}$, (3.42)
in mind, choosing
$\bar{s}_{k}=\binom{{x}}{{k+z}},\quad\bar{\sigma}_{k}=\binom{{k+x}}{{k+z}}$, (3.43)
in (3.41) yields
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{k+x}}{{k+z}}}=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{{k+z}}2^{n-k}}.$ (3.44)
for $n$ a non-negative integer and $x$ and $z$ complex numbers.
Proposition 3.5.
If $j$ and $n$ are non-negative integers, then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{n-j}{k-j}t_{n-k}}=\tau_{n-j}.$ (3.45)
Proof.
Consider the identity
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{j}}=(-1)^{j}\delta_{nj},$ (3.46)
and use $s_{k}=\binom{k}{j}$ and $\sigma_{k}=(-1)^{j}\delta_{kj}$ in (2.1). ∎
4 Identities involving mixed binomial transform pairs
Theorem 4.1.
Let $n$ be a non-negative integer. If $\{(s_{k}),(\sigma_{k})\}$ is a binomial transform pair of the first kind and $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, is a binomial-transform pair of the second kind, then
$\sum_{k=0}^{n}{\binom{{n}}{k}s_{k}\bar{t}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}\bar{\tau}_{n-k}}.$ (4.1)
Proof.
Let $\mathcal{M}_{x}(x^{k})=\bar{t}_{k}$ and $\mathcal{L}_{y}(y^{k})=s_{k}$.
Act on the identity
$\sum_{k=0}^{n}{\binom{{n}}{k}y^{n-k}x^{k}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\left({1-y}\right)^{n-k}\left({1+x}\right)^{k}},$
with $\mathcal{M}_{x}$ and $\mathcal{L}_{y}$, in succession. ∎
Example 4.1.
Use of the binomial-transform pair of the first kind $\{(F_{k}),(-F_{k})\}$, and the binomial-transform pair of the second kind $\{(B_{k}),((-1)^{k}B_{k})\}$, $k=0,1,2,\ldots$, in (4.1) gives
$\sum_{k=0}^{n}{\binom{{n}}{k}F_{k}B_{n-k}}=0,$ (4.2)
for $n$ an even integer.
Similarly $\{(L_{k}),(-L_{k})\}$ and $\{(B_{k}),((-1)^{k}B_{k})\}$, $k=0,1,2,\ldots$, in (4.1) yields
$\sum_{k=0}^{n}{\binom{{n}}{k}L_{k}B_{n-k}}=0,$ (4.3)
for $n$ an odd integer.
Example 4.2.
Use of the following binomial-transform pair (see Example 2.4):
$s_{k}=H_{k+m},\quad\sigma_{k}=\frac{{\delta_{k0}(1+H_{m})-1}}{{k+\delta_{k0}}}\binom{{k+m}}{m}^{-1},$
in (4.1) leads to the following identity
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k+m}\bar{t}_{n-k}}=H_{m}\bar{\tau}_{n}-\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\frac{1}{k}\,\bar{\tau}_{n-k}},$ (4.4)
which is valid for every binomial-transform pair of the second kind $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$ and every complex number $m$ that is not a negative integer.
In particular, the following polynomial identity holds:
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k+m}t^{n-k}}=(1+t)^{n}H_{m}-\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\frac{(1+t)^{n-k}}{k}},$
and can be recast as stated in proposition 4.1 by writing $1/t$ for $t$.
Proposition 4.1.
If $n$ is a non-negative integer, $m$ is a complex number that is not a negative integer and $t$ is a complex variable, then
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k+m}t^{k}}=\left({1+t}\right)^{n}H_{m}+\sum_{k=1}^{n}{(-1)^{k-1}\binom{{n}}{k}\binom{{k+m}}{m}^{-1}\frac{1}{k}\,t^{k}\left({1+t}\right)^{n-k}}.$ (4.5)
In particular, evaluation at $m=0$ and m=1/2m=-1/2, respectively, gives
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k}t^{k}}=\sum_{k=1}^{n}{(-1)^{k-1}\binom{{n}}{k}\frac{1}{k}\,t^{k}\left({1+t}\right)^{n-k}}$ (4.6)
and
$\sum_{k=0}^{n}{\binom{{n}}{k}O_{k}t^{k}}=\sum_{k=1}^{n}{(-1)^{k-1}\binom{{n}}{k}2^{2k-1}\binom{2k}{k}^{-1}\frac{1}{k}\,t^{k}\left({1+t}\right)^{n-k}},$ (4.7)
with the special values
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{k}}=\sum_{k=1}^{n}{(-1)^{k-1}2^{n-k}\binom{{n}}{k}\frac{1}{k}}$ (4.8)
and
$\sum_{k=0}^{n}{\binom{{n}}{k}O_{k}}=\sum_{k=1}^{n}{(-1)^{k-1}2^{n+k-1}\binom{{n}}{k}\binom{2k}{k}^{-1}\frac{1}{k}}.$ (4.9)
In deriving (4.7), we used (3.17) and (3.19).
Corollary 4.2.
Let $n$ be a non-negative integer. If $\{(s_{k}),(\sigma_{k})\}$ is a binomial transform pair of the first kind and $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, is a binomial-transform pair of the second kind, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\sigma_{k}\bar{t}_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}\bar{\tau}_{n-k}}.$ (4.10)
5 Binomial transform of products
Consider a binomial-transform pair of the second kind, say $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$ and an arbitrary sequence of complex numbers $(s_{k})$, $k=0,1,2,\ldots$. In this section we evaluate the binomial transform of the sequence $(s_{k}\bar{t}_{k})$ in terms of $(s_{k})$ and $(\bar{\tau}_{k})$, by establishing the result stated in the next theorem.
Theorem 5.1.
Let $n$ be a non-negative integer. If $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, is a binomial-transform pair of the second kind and $(s_{k})$, $k=0,1,2,\ldots$, is a sequence of complex numbers, then
$\sum_{k=0}^{n}{\binom{{n}}{k}s_{k}\bar{t}_{k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\bar{\tau}_{k}\sum_{j=0}^{n-k}{(-1)^{j}\binom{{n-k}}{j}s_{k+j}}}.$ (5.1)
Proof.
The elementary identity
$\sum_{k=0}^{n}{\binom{{n}}{k}x^{k}y^{k}}=\sum_{k=0}^{n}{\binom{{n}}{k}\left({1+y}\right)^{k}\sum_{j=0}^{n-k}{(-1)^{j}\binom{{n-k}}{j}x^{k+j}}}$ (5.2)
can be checked using the binomial theorem. Operating on both sides of (5.2) with $\mathcal{M}_{x}\mathcal{M}_{y}$ or $\mathcal{L}_{x}\mathcal{M}_{y}$ yields (5.1). ∎
Remark 5.1.
In terms of a binomial-transform pair of the first kind, $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, identity (5.1) reads
$\sum_{k=0}^{n}{\binom{{n}}{k}s_{k}t_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\tau_{k}\sum_{j=0}^{n-k}{\binom{{n-k}}{j}s_{k+j}}}.$ (5.3)
Theorem 5.2.
Let $n$ be a non-negative integer. If $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, is a binomial-transform pair of the second kind and $(s_{k})$, $k=0,1,2,\ldots$, is a sequence of complex numbers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}s_{k}\bar{t}_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}\bar{\tau}_{k}\sum_{j=0}^{n-k}{\binom{{n-k}}{j}s_{k+j}}}.$ (5.4)
Proof.
Replace $s_{k}$ with $(-1)^{k}s_{k}$ in (5.1) since $(s_{k})$ is an arbitrary sequence. ∎
Remark 5.2.
Written for a binomial-transform pair of the first kind, identity (5.4) becomes
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}s_{k}t_{k}}=\sum_{k=0}^{n}{\binom{n}{k}\tau_{k}\sum_{j=0}^{n-k}{(-1)^{j}\binom{{n-k}}{j}s_{k+j}}}.$ (5.5)
Remark 5.3.
Boyadzhiev [7] studied binomial transform of products. His results connect binomial-transform pairs and are expressed in terms of the forward difference operator.
Example 5.1.
Setting $s_{k}=G_{k}$ in (5.1) and making use of (2.8) gives
$\sum_{k=0}^{n}{\binom{{n}}{k}G_{k}\bar{t}_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{\tau}_{k}\left({G_{0}L_{n-2k}-G_{n-2k}}\right)}.$ (5.6)
The choice $s_{k}=G_{k}$ in (5.4) gives
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}G_{k}\bar{t}_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{\tau}_{k}G_{2n-k}};$ (5.7)
since [30, Equation (49)]:
$\sum_{j=0}^{n-k}{\binom{{n-k}}{j}G_{j+k}}=G_{2n-k}.$
Identity (5.7) corresponds to, but is slightly more general than, Boyadzhiev’s identity [7, Equation (1.11)].
6 Symmetry properties and generalizations
In this section we prove two symmetry properties and, based upon them, derive two generalizations of (2.1).
Theorem 6.1.
Let $\{(s_{k}),(\sigma_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. If $m$ and $n$ are non-negative integers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k+m}}=\sum_{k=0}^{m}{(-1)^{k}\binom{{m}}{k}\sigma_{k+n}}.$ (6.1)
Proof.
Let $\mathcal{L}_{x}(x^{k})=s_{k}$ for every complex number $x$ and every non-negative integer $k$. Operate on both sides of the following variation on the binomial theorem with $\mathcal{L}_{x}$:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}x^{k+m}}=\sum_{k=0}^{m}{(-1)^{k}\binom{{m}}{k}\left({1-x}\right)^{{}_{k+n}}}.$
Remark 6.1.
The version of (6.1) involving a binomial-transform pair of the second kind, namely,
$\sum_{k=0}^{n}{\binom{{n}}{k}\bar{s}_{k+m}}=\sum_{k=0}^{m}{(-1)^{k+m}\binom{{m}}{k}\bar{\sigma}_{k+n}},$ (6.2)
corresponds to Chen [9, Proof of Theorem 3.1, generalized in Theorem 3.2]. See also Gould and Quaintance [13] for an alternative proof of (6.2).
We now present a generalization of (2.1).
Theorem 6.2.
Let $m$, $n$ and $r$ be non-negative integers. If $\{(s_{k}),(\sigma_{k})\}$ and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the first kind, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{n-k+m}\sum_{q=0}^{r}{(-1)^{q}\binom{{r}}{q}\tau_{k+q}}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}t_{n-k+r}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}}}.$ (6.3)
Proof.
From (6.1) it is clear that $s_{k+m}$ and $\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}}$ are a binomial-transform pair of the first kind. That is, if $\{(s_{k}),(\sigma_{k})\}$, $k=0,1,2,\ldots$, is a binomial-transform pair of the first kind, then so is $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\ldots$, where
$a_{k}=s_{k+m}\text{ and }\alpha_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}}.$ (6.4)
Similarly, $\{(b_{k}),(\beta_{k})\}$, $k=0,1,2,\ldots$ is a binomial transform pair of the first kind, where
$b_{k}=t_{k+r}\text{ and }\beta_{k}=\sum_{q=0}^{r}{(-1)^{q}\binom{{r}}{q}\tau_{k+q}}.$ (6.5)
Now write (2.1) as
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}a_{n-k}b_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\alpha_{k}\beta_{n-k}},$ (6.6)
and substitute (6.4) and (6.5) to obtain
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}s_{k+m}t_{n-k+r}}$
$=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}\sum_{q=0}^{r}{(-1)^{q}\binom{{r}}{q}\tau_{n-k+q}}}},$ (6.7)
which can also be written as (6.2). ∎
We illustrate Theorem 6.2 with the result stated in Proposition 6.1.
Proposition 6.1.
If $j$, $m$, $r$ and $u$ are complex numbers and $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{n-k+m}}{u}\binom{{r}}{{j-k}}}=\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{n-k+r}}{j}\binom{{m}}{{u-k}}}$. (6.8)
Proof.
Use
$s_{k}=\binom{{k}}{u},\quad\sigma_{k}=(-1)^{u}\delta_{ku},$
and
$t_{k}=\binom{{k}}{j},\quad\tau_{k}=(-1)^{j}\delta_{kj},$
in (6.2). ∎
Theorem 6.3.
Let $m$ and $r$ be non-negative integers. If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, are a binomial-transform pair of the first kind, so are the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{k+p+m}}\text{ and }\beta_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p+r}}.$ (6.9)
Thus,
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{k+p+m}}}=\sum_{k=0}^{m}{(-1)^{k}\binom{m}{k}\sigma_{n+k+r}}.$ (6.10)
Proof.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. From (6.1), for an arbitrary non-negative integer $m$, we recognize
$s_{k+m}\text{ and }\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}},\quad k=0,1,2,\ldots,$ (6.11)
as a binomial-transform pair of the first kind.
Let $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. By the same token, for an arbitrary non-negative integer $r$,
$t_{k+r}\text{ and }\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}\tau_{k+p}},\quad k=0,1,2,\ldots,$ (6.12)
are a binomial-transform pair of the first kind.
In (6.11), let
$\tau_{k}=s_{k+m}\text{ and }t_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}},\quad k=0,1,2,\ldots$ (6.13)
Thus, using (6.13), the binomial transform pair given in (6.12) consists of
$t_{k+r}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p+r}}$ (6.14)
and
$\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}\tau_{k+p}}=\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{k+p+m}}.$ (6.15)
The proof is complete. ∎
In the next theorem, we present another generalization of (2.1).
Theorem 6.4.
Let $m$, $n$, $r$, $u$ and $v$ be non-negative integers. If $\{(s_{k}),(\sigma_{k})\}$ and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the first kind, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{n-k+p+m}}\sum_{j=0}^{u}{(-1)^{j}\binom{{u}}{j}t_{k+j+v}}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p+r}}\sum_{j=0}^{v}{(-1)^{j}\binom{{v}}{j}\tau_{n-k+j+u}}}.$ (6.16)
Proof.
In (6.6) substitute
$a_{k}=\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{k+p+m}},\quad\alpha_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p+r}},$
and
$b_{k}=\sum_{j=0}^{u}{(-1)^{j}\binom{{u}}{j}t_{k+j+v}},\quad\beta_{k}=\sum_{j=0}^{v}{(-1)^{j}\binom{{v}}{j}\tau_{k+j+u}}.$
7 Relations between binomial transforms
We derive various relations between binomial-transform pairs of both kinds. We develop relations which allow one to construct new binomial-transform pairs from known ones. Theorems 7.1 to 7.3 are derived from the binomial theorem.
Theorem 7.1.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=r,r+1,r+2,\dots$, where
$a_{k}=\binom{{k}}{r}t_{k-r}\text{ and }\alpha_{k}=(-1)^{r}\binom{{k}}{r}\tau_{k-r}.$
In particular, $(kt_{k-1})$ and $(-k\tau_{k-1})$, $k=1,2,3,\ldots$, constitute a binomial-transform pair of the first kind.
Proof.
Differentiate the binomial theorem
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}x^{k}}=\left({1-x}\right)^{n},$ (7.1)
$r$ times with respect to $x$ to obtain
$\sum_{k=r}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{r}x^{k-r}}=(-1)^{r}\binom{{n}}{r}\left({1-x}\right)^{n-r}$
and operate on both sides with $\mathcal{L}_{x}$. ∎
Theorem 7.2.
Let $n$ and $r$ be non-negative integers. If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=r,r+1,r+2,\dots$, where
$a_{k}=\binom{{k}}{r}t_{k}\text{ and }\alpha_{k}=\binom{{k}}{r}\sum_{j=0}^{r}{(-1)^{j}\binom{{r}}{j}\tau_{k-j}}.$
Thus,
$\sum_{k=r}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{r}t_{k}}=\binom{n}{r}\sum_{k=0}^{r}{(-1)^{k}\binom{{r}}{k}\tau_{n-k}}.$ (7.2)
Proof.
Write the binomial theorem as
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n-r}}{k}x^{k}}=\left({1-x}\right)^{n-r},\quad n\geq r$
and use the fact that
$\binom{{n-r}}{k}\binom{{n}}{r}=\binom{{n-k}}{r}\binom{{n}}{k}$
to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{r}x^{-k}}=(-1)^{n}\binom{{n}}{r}x^{-n}\left({1-x}\right)^{n-r}.$
Now write $1/x$ for $x$, simplify and replace $x$ with 1x1-x to get
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{r}\left({1-x}\right)^{k}}=\binom{n}{r}\sum_{k=0}^{r}{(-1)^{k}\binom{{r}}{k}x^{n-k}},$ (7.3)
from which (7.2) follows. ∎
Theorem 7.3.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=\frac{{t_{k+1}}}{{k+1}}\text{ and }\alpha_{k}=\frac{{\tau_{0}-\tau_{k+1}}}{{k+1}}.$ (7.4)
Thus,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{t_{k+1}}}{{k+1}}}=\frac{{\tau_{0}-\tau_{n+1}}}{{n+1}}.$ (7.5)
Proof.
Integrate the binomial theorem (7.1) from 0 to $y$ and replace $y$ with $x$. This gives
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{\left({1-x}\right)^{k+1}}}{{k+1}}}=\frac{{1-x^{n+1}}}{{n+1}}.$
Now act on both sides with $\mathcal{L}_{x}$ to obtain (7.5). ∎
Remark 7.1.
Each repeated application of Theorem 7.3 produces a new binomial-transform pair of the first kind. Given the binomial-transform pair $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$ and starting with (7.4) as
$t_{k}^{(1)}=\frac{{t_{k+1}}}{{k+1}}\text{ and }\tau_{k}^{(1)}=\frac{{\tau_{0}-\tau_{k+1}}}{{k+1}},$
we can derive a new binomial transform pair, namely $\{t_{k}^{(2)},\tau_{k}^{(2)}\}$, for $k=0,1,2,\ldots$, where
$t_{k}^{(2)}=\frac{{t_{k+1}^{(1)}}}{{k+1}}=\frac{t_{k+2}}{(k+1)(k+2)},$ (7.6)
and
$\tau_{k}^{(2)}=\frac{{\tau_{0}^{(1)}-\tau_{k+1}^{(1)}}}{{k+1}}=\frac{{\tau_{0}-\tau_{1}}}{{k+1}}-\frac{{\left({\tau_{0}-\tau_{k+2}}\right)}}{{(k+1)(k+2)}}.$ (7.7)
From (7.6) and (7.7) we can develop $\{t_{k}^{(3)},\tau_{k}^{(3)}\}$, for $k=0,1,2,\ldots$, where
$t_{k}^{(3)}=\frac{{t_{k+1}^{(2)}}}{{k+1}}=\frac{{t_{k+3}}}{{(k+1)(k+2)(k+3)}}$ (7.8)
and
$\tau_{k}^{(3)}$ $=\frac{{\tau_{0}^{(2)}-\tau_{k+1}^{(2)}}}{{k+1}}$
$=\frac{{\tau_{0}-\tau_{1}}}{{k+1}}-\frac{{\tau_{0}-\tau_{2}}}{{2(k+1)}}-\frac{{\tau_{0}-\tau_{1}}}{{(k+1)(k+2)}}+\frac{{\tau_{0}-\tau_{k+3}}}{{(k+1)(k+2)(k+3)}}.$ (7.9)
The consequence of 7.3 stated in Corollary 7.4 can be proved by induction.
Corollary 7.4.
Let $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. If $\tau_{0}=0=\tau_{1}=\cdots=\tau_{r-1}$, then $(t_{k}^{(r)})$ and $(\tau_{k}^{(r)})$ are a binomial transform pair of the first kind, where
$t_{k}^{(r)}=\frac{{t_{k+r}}}{{\left({k+1}\right)\left({k+2}\right)\cdots\left({k+r}\right)}}\text{ and }\tau_{k}^{(r)}=\frac{{(-1)^{r}\tau_{k+r}}}{{\left({k+1}\right)\left({k+2}\right)\cdots\left({k+r}\right)}},\quad r\in\mathbb{Z}^{+},$ (7.10)
or more compactly,
$t_{k}^{(r)}=\binom{{k+r}}{r}^{-1}t_{k+r}\text{ and }\tau_{k}^{(r)}=(-1)^{r}\binom{{k+r}}{r}^{-1}\tau_{k+r}.$ (7.11)
Example 7.1.
From the identity [15, Identity (6.17)]:
[…]
we identify the binomial-transform pair $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, where
[…]
Let $r$ be a positive integer. By (1.9), we have that if m>r1m>r-1, then $\tau_{0}=0=\tau_{1}=\tau_{2}=\cdots=\tau_{r-1}$. Using this fact in Corollary 7.4 gives the result stated in Proposition 7.1.
Proposition 7.1.
Let $m$ and $r$ be non-negative integers such that $m\geq r$. Then $(t_{k}^{(r)})$ and $(\tau_{k}^{(r)})$ are a binomial-transform pair of the first kind, where
[…]
Thus,
[…]
In particular,
[…]
Example 7.2.
Consider the following identity [5]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{m}}\binom{{s+k}}{k}=(-1)^{n}\binom{{n}}{m}\binom{{s+m}}{n}$, (7.17)
from which we identify the binomial-transform pair of the first kind $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, where
$t_{k}=(-1)^{k}\binom{{k}}{m}\binom{{s+m}}{k}\text{ and }\tau_{k}=\binom{{k}}{m}\binom{{s+k}}{k}$. (7.18)
Use of (7.18) in Corollary 7.4 produces the next result.
Proposition 7.2.
Let $s$ be a complex number that is not a negative integer. Let $m$ and $r$ be non-negative integers such that $m\geq r$. Then $(t_{k}^{(r)})$ and $(\tau_{k}^{(r)})$ are a binomial-transform pair of the first kind, where
$t_{k}^{(r)}=(-1)^{k}\binom{{k+r}}{m}\binom{{k+r}}{k}^{-1}\binom{{s+m}}{{k+r}}\text{ and }\tau_{k}^{(r)}=\binom{{k+r}}{m}\binom{{k+r}}{k}^{-1}\binom{{k+s+r}}{s}$. (7.19)
Thus,
$\sum_{k=0}^{n}{\binom{n}{k}\binom{{k+r}}{m}\binom{{k+r}}{k}^{-1}\binom{{s+m}}{{k+r}}}=\binom{{n+r}}{m}\binom{{n+r}}{n}^{-1}\binom{{n+s+r}}{s}$. (7.20)
In particular,
$\sum_{k=0}^{n}{\binom{n}{k}\binom{{s+r}}{{k+r}}}=\binom{{n+s+r}}{s}$. (7.21)
Theorem 7.5.
Let $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. Let $r$ be a non-negative integer. Then $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, is also a binomial-transform pair of the first kind, where
$t_{k}=\binom{{k+r}}{r}^{-1}a_{k+r}$ (7.22)
and
$\tau_{k}=(-1)^{r}\binom{{k+r}}{r}^{-1}\alpha_{k+r}-(-1)^{r}\binom{{k+r}}{r}^{-1}\sum_{j=0}^{r-1}{(-1)^{j}\binom{{k+r}}{j}a_{j}}.$ (7.23)
Proof.
Let $r$ be a non-negative integer. Let $(a_{k})$ be a sequence of complex numbers, where
[…]
where $x$ is a complex number. The inverse transform of $(a_{k})$ is computed as
$\alpha_{k}$ $=\sum_{j=0}^{k}{(-1)^{j}\binom{{k}}{j}a_{j}}$
$=\sum_{j=0}^{k}{(-1)^{j}\binom{{k}}{j}x^{j}}-\sum_{j=0}^{r-1}{(-1)^{j}\binom{{k}}{j}x^{j}}$
$=\left({1-x}\right)^{k}-\sum_{j=0}^{r-1}{(-1)^{j}\binom{{k}}{j}x^{j}}.$
By Corollary 7.4, the sequences $(a_{k}^{{}^{\prime}})$ and $(\alpha_{k}^{{}^{\prime}})$ where
$a_{k}^{{}^{\prime}}=\binom{k+r}{r}^{-1}a_{k+r}\text{ and }\alpha_{k}^{{}^{\prime}}=(-1)^{r}\binom{k+r}{r}^{-1}\alpha_{k+r}$ (7.25)
are a binomial-transform pair of the first kind. Thus,
$\sum_{j=0}^{k}{(-1)^{j}\binom{k}{j}\binom{{j+r}}{r}^{-1}x^{j+r}}$
$=(-1)^{r}\binom{{j+r}}{r}^{-1}\left({1-x}\right)^{k+r}-(-1)^{r}\binom{{k+r}}{r}^{-1}\sum_{j=0}^{r-1}{(-1)^{j}\binom{{k+r}}{j}x^{j}}.$ (7.26)
The statement of the theorem now follows upon application of Lemma 2.1 to (7). ∎
Remark 7.2.
Identity (7) is entry (4.13) in Gould’s book [14, p.47].
Example 7.3.
If we consider the anti-self-inverse sequence $(a_{k})$ and $(\alpha_{k})$, where $a_{k}=F_{k}=-\alpha_{k}$, we see on account of Theorem 7.5 that $t_{k}$ and $\tau_{k}$ are a binomial-transform pair of the first kind, where
$t_{k}=\binom{{k+r}}{r}^{-1}$ (7.27)
and
$\tau_{k}=(-1)^{r-1}\binom{{k+r}}{r}^{-1}F_{k+r}-(-1)^{r}\binom{{k+r}}{r}^{-1}\sum_{j=0}^{r-1}{(-1)^{j}\binom{{k+r}}{r}F_{j}}.$ (7.28)
Thus,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+r}}{r}^{-1}F_{k+r}}$
$\qquad=(-1)^{r-1}\binom{{n+r}}{r}^{-1}F_{n+r}-(-1)^{r}\binom{{n+r}}{r}^{-1}\sum_{j=0}^{r-1}{(-1)^{j}\binom{{n+r}}{j}F_{j}};$ (7.29)
with the corresponding result for the Lucas sequence being
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+r}}{r}^{-1}L_{k+r}}$
$\qquad=(-1)^{r}\binom{{n+r}}{r}^{-1}L_{n+r}-(-1)^{r}\binom{{n+r}}{r}^{-1}\sum_{j=0}^{r-1}{(-1)^{j}\binom{{n+r}}{j}L_{j}};$ (7.30)
Proposition 7.3.
Let $n$ be a non-negative integer. If $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$ is a binomial-transform pair of the second kind, then
$\sum_{k=0}^{n}{\bar{\tau}_{k}}=\sum_{k=0}^{n}{\binom{{n+1}}{{k+1}}\bar{t}_{k}}.$ (7.31)
In particular, if $x$ is a complex variable, then
$\sum_{k=0}^{n}{(1+x)^{k}}=\sum_{k=0}^{n}{\binom{{n+1}}{{k+1}}x^{k}}.$ (7.32)
Proof.
Set $m=n$ and $p=0$ in (3.30). ∎
Remark 7.3.
Shifting the summation index, identity (7.31) can also be written as
$\sum_{k=1}^{n}{\bar{\tau}_{k-1}}=\sum_{k=1}^{n}{\binom{n}{k}\bar{t}_{k-1}};$ (7.33)
which for a binomial transform pair of the first kind $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, also means
$\sum_{k=1}^{n}{\tau_{k-1}}=\sum_{k=1}^{n}{(-1)^{k-1}\binom{n}{k}t_{k-1}},$ (7.34)
in view of Remark 12.1.
Theorem 7.6.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=-\sum_{j=1}^{k}{\tau_{j-1}}\text{ and }\alpha_{k}=(1-\delta_{k0})t_{k-1+\delta_{k0}}.$
Proof.
An immediate consequence of (7.34). ∎
Example 7.4.
Sequences $(L_{k+1}-1)$ and $(-L_{k-1})$, $k=1,2,\ldots$, are a binomial-transform pair of the first kind since, if we take $t_{k}=L_{k}=\tau_{k}$, we have
$\sum_{j=1}^{k}L_{j-1}=L_{k+1}-1.$
Similarly, $(F_{k+1}-1)$ and $(F_{k-1})$, $k=1,2,\ldots$, are a binomial-transform pair of the first kind.
Theorem 7.7.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=\sum_{j=1}^{k-1+\delta_{k0}}{t_{j-1}}\text{ and }\beta_{k}=\sum_{j=1}^{k-1+\delta_{k0}}{\tau_{j-1}};$
so that
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{j=1}^{k-1}{t_{j-1}}}=\sum_{k=1}^{n-1}{\tau_{k-1}},\quad n\in\mathbb{Z}^{+}.$ (7.35)
In particular, if $(t_{k})$, $k=0,1,2,\dots$, is a self-inverse sequence, so is the sequence $\left(\sum_{j=1}^{k-1+\delta_{k0}}{t_{j-1}}\right)$.
Proof.
With Theorem 7.6 in mind, let $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\dots$, be a binomial-transform pair of the first kind, where
$a_{k}=-\sum_{j=1}^{k}{\tau_{j-1}},\quad\alpha_{k}=\left({1-\delta_{k0}}\right)t_{k-1+\delta_{k0}},$ (7.36)
and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\dots$, is a binomial-transform pair of the first kind.
By the same Theorem 7.6, $\sum_{j=1}^{k}{\alpha_{j-1}}$ and $\left({1-\delta_{k0}}\right)a_{k-1+\delta_{k0}}$ are a binomial-transform pair of the first kind.
Using (7.36), we have
$-\sum_{j=1}^{k}{\alpha_{j-1}}$ $=-\sum_{j=1}^{k}{\left({1-\delta_{j1}}\right)t_{j-2+\delta_{j1}}}$
$=-\sum_{j=1}^{k}{t_{j-2+\delta_{j1}}}+\sum_{j=1}^{k}{\delta_{j1}t_{j-2+\delta_{j1}}}$
$=-t_{0}-\sum_{j=2}^{k}{t_{j-2}}+t_{0}$
$=-\sum_{j=2}^{k}{t_{j-2}}=-\sum_{j=1}^{k-1}{t_{j-1}}.$
Since the left side is an empty sum for $k=0$, in order to preserve equality, we therefore write
$\sum_{j=1}^{k}{\alpha_{j-1}}=-\sum_{j=1}^{k-1+\delta_{k0}}{t_{j-1}}.$
From (7.36), we also have
$\left({1-\delta_{k0}}\right)a_{k-1+\delta_{k0}}=-\left({1-\delta_{k0}}\right)\sum_{j=1}^{k-1+\delta_{k0}}{\tau_{j-1}}=-\sum_{j=1}^{k-1+\delta_{k0}}{\tau_{j-1}}.$
Theorem 7.8.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=\frac{1}{{\left({k+1}\right)\left({k+2}\right)}}\sum\limits_{j=0}^{k}{t_{j}}\text{ and }\alpha_{k}=\frac{1}{{\left({k+1}\right)\left({k+2}\right)}}\sum\limits_{j=0}^{k}{\tau_{j}}.$ (7.37)
In particular, if $(t_{k})$, $k=0,1,2,\dots$, is a self-inverse sequence, so is the sequence
$\left(\frac{1}{{\left({k+1}\right)\left({k+2}\right)}}\sum\limits_{j=0}^{k}{t_{j}}\right).$
Proof.
By shifting the index, identity (7.35) can be written as
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n+2}}{{k+2}}\sum_{j=0}^{k}{t_{j}}}=\sum_{k=0}^{n}{\tau_{k}},$
from which (7.37) follows since
$\binom{{n+2}}{{k+2}}=\frac{{\left({n+2}\right)\left({n+1}\right)}}{{\left({k+2}\right)\left({k+1}\right)}}\binom{{n}}{k}$.
Example 7.5.
The sequence
$\left(\frac{{L_{k+2}-1}}{(k+1)(k+2)}\right),\quad k=0,1,2,\ldots$
is a self-inverse sequence, since $(L_{k})$ is a self-inverse sequence and
$\sum_{k=0}^{n}L_{k}=L_{n+2}-1.$
Theorem 7.9.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\tau_{j}}\text{ and }\alpha_{k}=\frac{{t_{k}}}{{k+1}}.$ (7.38)
Proof.
Identity (7.31) written for a binomial-transform pair of the first kind is
$\sum_{k=0}^{n}{\tau_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n+1}}{{k+1}}t_{k}},$
which can also be written as
$\frac{1}{{n+1}}\sum_{k=0}^{n}{\tau_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{t_{k}}}{{k+1}}},$
and hence the theorem. ∎
Theorem 7.10.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=-\frac{1}{{k+1}}\sum\limits_{j=1}^{k}{\tau_{j-1}}\text{ and }\beta_{k}=\frac{1}{{k+1}}\sum\limits_{j=1}^{k}{t_{j-1}}.$ (7.39)
In particular, if $(t_{k})$, $k=0,1,2,\dots$, is a self-inverse sequence, then the sequence
$\left(\frac{1}{{k+1}}\sum\limits_{j=1}^{k}{t_{j-1}}\right),\quad k=0,1,2,\dots,$
is an anti-self-inverse sequence, and vice versa.
Proof.
Let $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\dots$, be a binomial-transform pair of the first kind, where
$a_{k}=-\sum_{j=1}^{k}{\tau_{j-1}},\quad\alpha_{k}=\left({1-\delta_{k0}}\right)t_{k-1+\delta_{k0}},$
and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind.
Then
$\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\alpha_{j}}=\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\left({1-\delta_{j0}}\right)t_{j-1+\delta_{j0}}}=\frac{1}{{k+1}}\sum\limits_{j=1}^{k}{t_{j-1}}$
and
$\frac{{a_{k}}}{{k+1}}=-\frac{1}{{k+1}}\sum\limits_{j=1}^{k}{\tau_{j-1}}$
are a binomial-transform pair of the first kind, by Theorem 7.9. ∎
Theorem 7.11.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\frac{{t_{j}}}{{j+1}}}\text{ and }\beta_{k}=\frac{1}{{\left({k+1}\right)^{2}}}\sum\limits_{j=0}^{k}{\tau_{j}}.$ (7.40)
Proof.
Let $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\dots$, be a binomial-transform pair of the first kind, where
$a_{k}=\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\tau_{k}},\quad\alpha_{k}=\frac{{t_{k}}}{{k+1}},$
and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\dots$, is a binomial-transform pair of the first kind.
Then the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\alpha_{j}}=\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\frac{{t_{j}}}{{j+1}}}$
and
$\beta_{k}=\frac{{a_{k}}}{{k+1}}=\frac{1}{{\left({k+1}\right)^{2}}}\sum\limits_{j=0}^{k}{\tau_{j}}$
are a binomial-transform pair of the first kind, by Theorem 7.9. ∎
Theorem 7.12.
Let $m$ be a non-negative integer. If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=\frac{1}{{k+1}}\sum_{j=0}^{k}{t_{j+m}}\text{ and }\beta_{k}=\frac{1}{{k+1}}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\tau_{k+p}}.$ (7.41)
Proof.
Let $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind, where (see Equation (6.4))
$a_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\tau_{k+p}}\text{ and }\alpha_{k}=t_{k+m}.$
By Theorem 7.9, sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\ldots$, given by
$b_{k}=\frac{1}{{k+1}}\sum_{j=0}^{k}{\alpha_{k}}=\frac{1}{{k+1}}\sum_{j=0}^{k}{t_{j+m}},\quad k=0,1,2,\ldots,$
and
$\beta_{k}=\frac{{a_{k}}}{{k+1}}=\frac{1}{{k+1}}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\tau_{k+p}},\quad k=0,1,2,\ldots,$
are also a binomial-transform pair of the first kind. ∎
Example 7.6.
By choosing $t_{k}=F_{k}$ and $\tau_{k}=-F_{k}$ in Theorem 7.12, we find a new binomial-transform pair of the first kind $\{(b_{k}),(\beta_{k})\}$, $k=0,1,2,\ldots$, where
$b_{k}=\frac{F_{m+k+2}-F_{m+1}}{k+1}\text{ and }\beta_{k}=\frac{(-1)^{k}F_{m-k}}{k+1},$ (7.42)
since [17, p.87]:
$\sum_{j=0}^{k}{F_{j+m}}=F_{k+m+2}-F_{m+1}$
and from (2.8):
$\sum_{p=0}^{m}{(-1)^{p}\binom{m}{p}F_{k+p}}=(-1)^{k-1}F_{m-k}.$
Theorem 7.13.
Let $m$ be a non-negative integer. If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=\frac{1}{{k+m+1}}\sum_{j=0}^{k+m}{\tau_{j}}\text{ and }\beta_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\frac{{t_{k+p}}}{{k+p+1}}}.$ (7.43)
Thus,
$\sum_{k=0}^{n}{\frac{{(-1)^{k}}}{{k+m+1}}\binom{{n}}{k}\sum_{j=0}^{k+m}{\tau_{j}}}=\sum_{k=0}^{m}{\frac{{(-1)^{k}}}{{k+n+1}}\binom{{m}}{k}t_{k+n}}.$ (7.44)
Proof.
Let $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind, where
$a_{k}=\frac{1}{{k+1}}\sum_{j=0}^{k}{\tau_{j}}\text{ and }\alpha_{k}=\frac{{t_{k}}}{{k+1}}.$
By (6.4), so also are $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\ldots$, where
$b_{k}=a_{k+m}=\frac{1}{{k+m+1}}\sum_{j=0}^{k+m}{\tau_{j}}$
and
$\beta_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{m}{p}\alpha_{k+p}}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\frac{{t_{k+p}}}{{k+p+1}}}.$
Theorem 7.14.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\dots$, where
$b_{k}=-\sum_{j=1}^{k}{\frac{{t_{j-1}}}{j}}\text{ and }\beta_{k}=\frac{1}{{k+\delta_{k0}}}\sum_{j=1}^{k}{\tau_{j-1}}.$ (7.45)
Proof.
Let
$a_{k}=\frac{1}{{k+1}}\sum_{j=0}^{k}{\tau_{j}}\text{ and }\alpha_{k}=\frac{{t_{k}}}{{k+1}},\quad k=0,1,2,\ldots,$
be a binomial-transform pair of the first kind. By Theorem 7.6, so also are
$\sum_{j=1}^{k}{\alpha_{j-1}}=-\sum_{j=1}^{k}{\frac{{t_{j-1}}}{j}}$
and
$\left({1-\delta_{k0}}\right)a_{k-1+\delta_{k0}}=\frac{1}{{k+\delta_{k0}}}\sum_{j=1}^{k}{\tau_{j-1}}.$
Theorem 7.15.
If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=-\sum_{j=1}^{k}{2^{-j+1}s_{j-1}}\text{ and }\alpha_{k}=\frac{{2^{-k+1}}}{k+\delta_{k0}}\sum_{j=1}^{k}{j\binom{{k}}{j}\sigma_{j-1}},\quad k=0,1,2,\ldots.$ (7.46)
Proof.
Consider the binomial-transform pair of the first kind version of (3.41), namely,
$\sum_{k=0}^{n}{\binom{{n}}{k}\sigma_{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{n-k}s_{k}}.$ (7.47)
We see that $(b_{k})$ and $(\beta_{k})$, $k=0,1,2\ldots$, are a binomial transform pair of the first kind, where
$b_{k}=2^{-k}\sum_{j=0}^{k}{\binom{k}{j}\sigma_{j}}\text{ and }\beta_{k}=2^{-k}s_{k}.$ (7.48)
By Theorem 7.6, $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2\ldots$, are also a binomial-transform pair of the first kind, where
$a_{k}=-\sum_{j=1}^{k}{\beta_{j-1}}\text{ and }\alpha_{k}=\left({1-\delta_{k0}}\right)b_{k-1+\delta_{k0}}.$
Theorem 7.16.
If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=\frac{1}{{k+1}}\sum_{j=0}^{k}{2^{-j}s_{j}}\text{ and }\alpha_{k}=\frac{{2^{-k}}}{{k+1}}\sum_{j=0}^{k}{\binom{{k}}{j}\sigma_{j}}.$ (7.49)
Proof.
Let $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\ldots$, be the binomial-transform pair of sequences defined in (7.48). By Theorem 7.9, $(a_{k})$ and $(\alpha_{k})$ are also a binomial-transform pair of the first kind, where
$a_{k}=\frac{1}{{k+1}}\sum_{j=0}^{k}{\beta_{j}}=\frac{1}{{k+1}}\sum_{j=0}^{k}{2^{-j}s_{j}}$
and
$\alpha_{k}=\frac{{a_{k}}}{{k+1}}=\frac{{2^{-k}}}{{k+1}}\sum_{j=0}^{k}{\binom{{k}}{j}\sigma_{j}}.$
Example 7.7.
By choosing $s_{k}=1$ and $\sigma_{k}=\delta_{k0}$ in (7.49), we find that
$\left(\frac{2-2^{-k}}{k+1}\right)\text{ and }\left(\frac{2^{-k}}{k+1}\right),\quad k=0,1,2,\ldots,$
are a binomial-transform pair of the first kind.
Theorem 7.17.
If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=\frac{{2^{-k}}}{{k+1}}\sum_{j=0}^{k}{\sigma_{j}}\text{ and }\alpha_{k}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}\frac{{s_{j}}}{{j+1}}}.$ (7.50)
Proof.
Using Theorem 7.9, consider the binomial-transform pair of the first kind $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\ldots$, where
$b_{k}=\frac{1}{{k+1}}\sum\limits_{j=0}^{k}{\sigma_{j}}\text{ and }\beta_{k}=\frac{{s_{k}}}{{k+1}}.$
By (7.48),
$a_{k}=2^{-k}b_{k}\text{ and }\alpha_{k}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}\beta_{j}}$
are also a binomial-transform pair of the first kind. ∎
Theorem 7.18.
If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}2^{-j}s_{j}}\text{ and }\alpha_{k}=2^{-2k}\sum_{j=0}^{k}{\binom{{k}}{j}\sigma_{j}}.$ (7.51)
Proof.
With (7.48) in mind, consider the binomial-transform pair of the first kind $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\ldots$, where
$b_{k}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}\sigma_{j}}\text{ and }\beta_{k}=2^{-k}s_{k}.$
By the same (7.48), $a_{k}$ and $\alpha_{k}$ are also a binomial-transform pair of the first kind, where
$a_{k}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}\beta_{j}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}2^{-j}s_{j}}},$
and
$\alpha_{k}=2^{-k}b_{k}=2^{-2k}\sum_{j=0}^{k}{\binom{{k}}{j}\sigma_{j}}.$
Example 7.8.
By choosing $s_{k}=1$ and $\sigma_{k}=\delta_{k0}$ in (7.51), we obtain
$\left(2^{-2k}3^{k}\right)\text{ and }\left(2^{-2k}\right),\quad k=0,1,2,\ldots,$
as a binomial-transform pair of the first kind.
Example 7.9.
By choosing $s_{k}=(1+k)^{-1}=\sigma_{k}$ in (7.51), we find that the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
$a_{k}=\frac{{2^{1-k}-2^{-2k}}}{{k+1}}\text{ and }\alpha_{k}=\frac{{2^{-2k}3^{k+1}-2^{1-k}}}{{k+1}},$
are a binomial-transform pair of the first kind.
Theorem 7.19.
If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=-2^{-k}\sum_{j=1}^{k}{\sigma_{j-1}}\text{ and }\alpha_{k}=2^{-k}\sum_{j=1}^{k}{\binom{{k}}{j}s_{j-1}}.$ (7.52)
Proof.
Let $(b_{k})$ and $(\beta_{k})$ be the binomial-transform pair of the first kind given by
$b_{k}=-\sum_{j=1}^{k}{\sigma_{j-1}}\text{ and }\beta_{k}=(1-\delta_{k0})s_{k-1+\delta_{k0}},$
where $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, are a binomial transform-pair of the first kind. By (7.48), it is also true that
$a_{k}=2^{-k}b_{k}\text{ and }\alpha_{k}=2^{-k}\sum_{j=0}^{k}{\binom{k}{j}}\beta_{j},$
are also a binomial-transform pair of the first kind. ∎
Theorem 7.20.
Let $m$ be a non-negative integer. If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=2^{-k}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{p+k}}\text{ and }\alpha_{k}=2^{-k}\sum_{p=0}^{k}{\binom{{k}}{p}s_{p+m}}.$ (7.53)
Proof.
Recall (see (6.4)) that $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\ldots$, are a binomial-transform pair of the first kind, where
$b_{k}=\sum_{k=0}^{m}{(-1)^{p}\binom{m}{p}\sigma_{k+p}}\text{ and }\beta_{k}=s_{k+m}.$
By (7.48), it is also a fact that
$a_{k}=2^{-k}b_{k}\text{ and }\alpha_{k}=2^{-k}\sum_{j=0}^{k}{\binom{k}{j}\beta_{j}}$
are a binomial-transform pair of the first kind. ∎
Example 7.10.
Choosing $s_{k}=L_{k}=\sigma_{k}$ in (7.53) gives the binomial-transform pair of the first kind $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
$a_{k}=(-1)^{k}2^{-k}L_{m-k}\text{ and }\alpha_{k}=2^{-k}L_{m+2k},$
since (Equation (2.8)):
$\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}L_{p+k}}=(-1)^{k}L_{m-k}$
and [30, Equation (49)]
$\sum_{p=0}^{k}{\binom{{k}}{p}L_{p+m}}=L_{m+2k}.$
Theorem 7.21.
Let $m$ be a non-negative integer. If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=2^{-k}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}2^{-p}\sigma_{k+p}}\text{ and }\alpha_{k}=2^{-k-m}\sum_{j=0}^{k+m}{\binom{{k+m}}{j}s_{j}}.$ (7.54)
Proof.
Let $(b_{k})$ and $(\beta_{k})$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind (see (7.48)), where
$b_{k}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}s_{j}}\text{ and }\beta_{k}=2^{-k}\sigma_{k},$
and $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, are a binomial-transform pair of the first kind. By the same (7.48), $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, constitute a binomial-transform pair of the first kind, where
$a_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\beta_{k+p}}=2^{-k}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}2^{-p}\sigma_{k+p}}$
and
$\alpha_{k}=b_{k+m}=2^{-k-m}\sum_{j=0}^{k+m}{\binom{{k+m}}{j}s_{j}}.$
Theorem 7.22.
Let $m$ be a non-negative integer. If $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\dots$, where
$a_{k}=-\sum_{j=1}^{k}{s_{m+j-1}}\text{ and }\alpha_{k}=\left({1-\delta_{k0}}\right)\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k-1+\delta_{k0}+p}}.$ (7.55)
Proof.
Consider the binomial transform pair of the first kind (Equation (6.4)), $\{(a_{k}),(\alpha_{k})\}$, $k=0,1,2,\ldots$, where
$b_{k}=\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\sigma_{k+p}}\text{ and }\beta_{k}=s_{k+m}.$
By Theorem 7.6, $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, are also a binomial-transform pair of the first kind, where
$a_{k}=-\sum_{j=1}^{k}{\beta_{j-1}}\text{ and }\alpha_{k}=(1-\delta_{k0})b_{k-1+\delta_{k0}}.$
Theorem 7.23.
Let $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. Then $\left(kt_{k}\right)$ and $\left(k(\tau_{k}-\tau_{k-1})\right)$ are a binomial-transform pair of the first kind for $k=1,2,\ldots$.
Proof.
Set $j=1$ in (3.45) on page 3.45 and use the identity
$\binom{n}{k}=\frac{n}{k}\binom{n-1}{k-1}$
to obtain
$n\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}t_{k}}-\sum_{k=0}^{n}{(-1)^{k}k\binom{{n}}{k}t_{k}}=n\tau_{n-1};$
and hence
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}kt_{k}}=n\left({\tau_{n}-\tau_{n-1}}\right).$ (7.56)
Example 7.11.
If on the basis of the recurrence relation of Bernoulli numbers, namely,
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}(-1)^{k}B_{k}}=(-1)^{n}B_{n},$
and use $t_{k}=(-1)^{k}B_{k}=\tau_{k}$ in (7.56), we obtain, for $n$ a non-negative integer,
$\sum_{k=0}^{n}{\binom{n}{k}kB_{k}}=(-1)^{n}n\left({B_{n}+B_{n-1}}\right),$ (7.57)
that is
$\sum_{k=0}^{n}{\binom{n}{k}kB_{k}}=\begin{cases}nB_{n},&\text{if $n$ is even};\\ -nB_{n-1},&\text{if $n>1$ is odd};\\ -1/2,&\text{if $n=1$}.\end{cases}$ (7.58)
Corollary 7.24.
Let $\{(s_{k}),(\sigma_{k})\}$, $k$ an integer, be a binomial-transform pair of the first kind. Let $S_{0}(k)=\sigma_{k}$ and $S_{m}(k)=k(S_{m-1}(k)-S_{m-1}(k-1))$ for every positive integer $m$. Then $S_{m}(k)$ and $k^{m}s_{k}$ are a binomial-transform pair of the first kind; that is
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}k^{m}s_{k}}=S_{m}(n),\quad m=0,1,2,\ldots.$ (7.59)
In particular,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}k^{2}s_{k}}=n^{2}\left({\sigma_{n}-\sigma_{n-1}}\right)-n(n-1)\left({\sigma_{n-1}-\sigma_{n-2}}\right),$ (7.60)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}k^{3}s_{k}}$ $=n^{3}\left({\sigma_{n}-\sigma_{n-1}}\right)-n\left({n-1}\right)\left({2n-1}\right)\left({\sigma_{n-1}-\sigma_{n-2}}\right)$
$\qquad+n\left({n-1}\right)\left({n-2}\right)\left({\sigma_{n-2}-\sigma_{n-3}}\right).$ (7.61)
Theorem 7.25.
If $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\dots$, are a binomial-transform pair of the first kind, then so are the sequences $(a_{k})$ and $(\alpha_{k})$, $k=1,2,3,\dots$, where
$a_{k}=\sum_{j=1}^{k}{t_{j}}\text{ and }\alpha_{k}=\tau_{k}-\tau_{k-1}.$
We also have that $(b_{k})$ and $(\beta_{k})$, $k=1,2,3,\dots$, are a binomial-transform pair of the first kind, where
$b_{k}=(-1)^{k}\sum_{j=1}^{k}{(-1)^{j}t_{j}}\text{ and }\beta_{k}=2^{k}\sum_{j=1}^{k}{2^{-j}\left({\tau_{j}-\tau_{j-1}}\right)}.$
Thus,
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{j=1}^{k}{t_{j}}}=\tau_{n}-\tau_{n-1}$ (7.62)
and
$\sum_{k=1}^{n}{\binom{{n}}{k}\sum_{j=1}^{k}{(-1)^{j}t_{j}}}=2^{n}\sum_{k=1}^{n}{2^{-k}\left({\tau_{k}-\tau_{k-1}}\right)}.$ (7.63)
Proof.
The following identities are known ($s=0$ in [4, Cor. 2, 4]):
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{k=1}^{n}{t_{k}}}=-\sum_{k=0}^{n-1}{(-1)^{k}\binom{{n-1}}{k}t_{k+1}},$ (7.64)
and
$\sum_{k=1}^{n}{\binom{{n}}{k}\sum_{k=1}^{n}{t_{k}}}=2^{n}\sum_{k=1}^{n}{\frac{1}{{2^{k}}}\sum_{j=0}^{k-1}{\binom{{k-1}}{j}t_{j+1}}}.$ (7.65)
Write (7.64) as
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\sum_{k=1}^{n}{t_{k}}}=\sum_{k=1}^{n}{(-1)^{k}\binom{{n-1}}{{k-1}}t_{k}}=\sum_{k=0}^{n}{(-1)^{k}\frac{k}{n}\binom{{n}}{k}t_{k}}$
and use (7.56) to obtain (7.62). The proof of (7.63) is similar. ∎
8 Self-inverse sequences
In this section, we discuss how to construct invariant or self-inverse sequences and anti-self-inverse sequences from given binomial transform pairs. The results derived here complement those found by previous researchers [11, 26, 31].
Theorem 8.1.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. Then, for $k=0,1,2,\ldots$, the sequence $(a_{k})$ given by $a_{k}=s_{k}+\sigma_{k}$ is a self-inverse sequence while the sequence $(b_{k})$ where $b_{k}=s_{k}-\sigma_{k}$ is an anti-self-inverse sequence.
Proof.
We have
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({s_{k}+\sigma_{k}}\right)}$ $=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}}+\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}}$
$=\sigma_{n}+s_{n},$
as claimed. Similarly,
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({s_{k}-\sigma_{k}}\right)}=\sigma_{n}-s_{n}=-(s_{n}-\sigma_{n}).$
In particular, for $k=0,1,2,\ldots$, the sequence $(a_{k})$ given by $a_{k}=x^{k}+(1-x)^{k}$ is a self-inverse sequence while the sequence $(b_{k})$ where $b_{k}=x^{k}-(1-x)^{k}$ is an anti-self-inverse sequence for every complex number $x$.
Theorem 8.2.
If $(t_{k})$, $k=0,1,2,\ldots,$ is an anti-self-inverse sequence, then the sequence $(a_{k})$, where
$a_{k}=\frac{t_{k+1}}{k+1}$ (8.1)
is a self inverse sequence.
Proof.
Clearly, $t_{0}=0=\tau_{0}$ for an anti-self-inverse sequence. Since here $(t_{k})$ is anti-self-inverse by hypothesis, we have $\tau_{k}=-t_{k}$ and hence (8.1) when we invoke Theorems 7.3 and 8.1. ∎
Remark 8.1.
Theorem 8.2 is known; see, for example, Wang [31].
Theorem 8.3.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. Then, for $k=0,1,2,\ldots$, the sequence $(a_{k})$ given by
$a_{k}=\frac{1}{k+1}\left(s_{k+1}-\sigma_{k+1}\right)$ (8.2)
is a self-inverse sequence.
Proof.
This is a consequence of Theorems 8.1 and 8.2. ∎
Theorem 8.4.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. Then, for $k=0,1,2,\ldots$, the sequence $(a_{k})$ given by
$a_{k}=H_{k-1+\delta_{k0}}$ (8.3)
is a self-inverse sequence while the sequence $(b_{k})$, where
$b_{k}=\frac{{kH_{k}+1-\delta_{k0}}}{{k+\delta_{k0}}}$ (8.4)
is an anti-self-inverse sequence.
Proof.
Since
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}H_{k}}=\begin{cases}-1/n,&\text{$n\neq 0$};\\ 0,&\text{$n=0$};\end{cases}$
we have that, for $k=0,1,2,\ldots$, sequences $(s_{k})$ and $(\sigma_{k})$ are a binomial-transform pair of the first kind, where
$s_{k}=H_{k}\text{ and }\sigma_{k}=-\frac{1-\delta_{k0}}{k+\delta_{k0}}.$ (8.5)
By Theorem 8.1, for $k=0,1,2,\ldots$, we have that $(a_{k})$ is a self-inverse sequence while $(b_{k})$ is an anti-self-inverse sequence, where
[…]
$=H_{k-1+\delta_{k0}};$
and
[…]
$=\frac{{1-\delta_{k0}}}{{k+\delta_{k0}}}\left({kH_{k}+1}\right).$
Remark 8.2.
The self-inverse sequence $(a_{k})$ and the anti-self-inverse sequence $(b_{k})$ are given more briefly by
$a_{k}=H_{k-1}\text{ and }b_{k}=\frac{\left({kH_{k}+1}\right)}{k},\quad k=1,2,3,\ldots$
Theorem 8.5.
For $k=1,2,3,\ldots$, the sequence $(s_{k})$ is an anti-self-inverse sequence, where
$s_{k}=(-1)^{k-1}kB_{k-1};$ (8.6)
so that
$\sum_{k=1}^{n}{\binom{{n}}{k}kB_{k-1}}=(-1)^{n-1}nB_{n-1}.$ (8.7)
Proof.
Follows from (7.57) and Theorem 8.1. ∎
Using Theorem 8.1, self-inverse sequences and anti-self-inverse sequences can be deduced from each of the binomial-transform pairs established in Section 7. We give some examples.
Theorem 8.6.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. Then, for $k=0,1,2,\ldots$, the sequence $(a_{k})$ given by
$a_{k}=\sum\limits_{j=1}^{k}{\sigma_{j-1}}-\left({1-\delta_{k0}}\right)s_{k-1+\delta_{k0}}$ (8.8)
is a self-inverse sequence while the sequence $(b_{k})$, where
$b_{k}=\sum\limits_{j=1}^{k}{\sigma_{j-1}}+\left({1-\delta_{k0}}\right)s_{k-1+\delta_{k0}}$ (8.9)
is an anti-self-inverse sequence.
Proof.
Follows from Theorem 7.6 and Theorem 8.1. ∎
Corollary 8.7.
If $(s_{k})$, $k=0,1,2,\ldots$, is a self-inverse sequence, then the sequence $(a_{k})$, where
$a_{k}=\sum_{j=1}^{k-1}s_{j-1},\quad k=1,2,3,\ldots,$ (8.10)
is a self-inverse sequence while the sequence $(b_{k})$, where
$b_{k}=\sum_{j=1}^{k-1}s_{j-1}+2s_{k-1},\quad k=1,2,3,\ldots,$ (8.11)
is an anti-self-inverse sequence. Similarly, if $(s_{k})$, $k=0,1,2,\ldots$, is an anti-self-inverse sequence, then the sequence $(b_{k})$, where
$b_{k}=\sum_{j=1}^{k-1}s_{j-1},\quad k=1,2,3,\ldots,$ (8.12)
is an anti-self-inverse sequence while the sequence $(a_{k})$, where
$a_{k}=\sum_{j=1}^{k-1}s_{j-1}+2s_{k-1},\quad k=1,2,3,\ldots,$ (8.13)
is a self-inverse sequence.
Corollary 8.7 is useful because it allows one to derive a new self-inverse sequence and a new anti-self-inverse sequence from a known self-inverse sequence or anti-self-inverse sequence. See also Corollary 8.11.
Example 8.1.
If we choose $s_{k}=H_{k-1+\delta_{k0}}$, $k=0,1,2,\ldots$ in Corollary 8.7 and use also (3.33), we find that $(a_{k})$ is a self-inverse sequence, where
$a_{k}=(k-2)(H_{k-2}-1),\quad k=2,3,4,\ldots,$ (8.14)
and $(b_{k})$ is an anti-self-inverse sequence, where
$b_{k}=k(H_{k-2}-1)+2,\quad k=2,3,4,\ldots$ (8.15)
Theorem 8.8.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. Then, for $k=1,2,3,\ldots$, the sequence $(a_{k})$ given by
$a_{k}=\sum_{j=1}^{k-1}{\left(s_{j-1}+\sigma_{j-1}\right)},$ (8.16)
is a self-inverse sequence while the sequence $(b_{k})$, where
$b_{k}=\sum_{j=1}^{k-1}{\left(s_{j-1}-\sigma_{j-1}\right)},$ (8.17)
is an anti-self-inverse sequence.
Proof.
A consequence of Theorem 7.7, using Theorem 8.1. ∎
Example 8.2.
The choice $s_{k}=x^{k}$ and $\sigma_{k}=(1-x)^{k}$ in Theorem 8.8 gives the self-inverse sequence $(a_{k})$, where
$a_{k}=\frac{1-x^{k-1}}{1-x}+\frac{1-(1-x)^{k-1}}{x},\quad k=1,2,3,\ldots,$ (8.18)
and the anti-self-inverse sequence $(b_{k})$, where
$b_{k}=\frac{1-x^{k-1}}{1-x}-\frac{1-(1-x)^{k-1}}{x},\quad k=1,2,3,\ldots,$ (8.19)
for every complex number $x$.
The next result comes from Theorem 7.8.
Theorem 8.9.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. Then, for $k=0,1,2,3,\ldots$, we have a self-inverse sequence $(a_{k})$ given by
$a_{k}=\frac{1}{{\left({k+1}\right)\left({k+2}\right)}}\sum_{j=0}^{k}{\left({s_{j}+\sigma_{j}}\right)},$ (8.20)
and an anti-self-inverse sequence $(b_{k})$ given by
$b_{k}=\frac{1}{{\left({k+1}\right)\left({k+2}\right)}}\sum_{j=0}^{k}{\left({s_{j}-\sigma_{j}}\right)}.$ (8.21)
Example 8.3.
Using (8.5) in Theorem 8.9, we establish, for $k=0,1,2,\ldots$, the self-inverse sequence $(a_{k})$, where
$a_{k}=\frac{k(H_{k}-1)}{(k+1)(k+2)},$ (8.22)
and the anti-self-inverse sequence $(b_{k})$, where
$b_{k}=\frac{(k+2)H_{k}-k}{(k+1)(k+2)}.$ (8.23)
The next result is from Theorem 7.10.
Theorem 8.10.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. Then, for $k=0,1,2,3,\ldots$, we have a self-inverse sequence $(a_{k})$ given by
$a_{k}=\frac{1}{k+1}\sum_{j=1}^{k}{\left({s_{j-1}-\sigma_{j-1}}\right)},$ (8.24)
and an anti-self-inverse sequence $(b_{k})$ given by
$b_{k}=\frac{1}{k+1}\sum_{j=1}^{k}{\left({s_{j-1}+\sigma_{j-1}}\right)}.$ (8.25)
Corollary 8.11.
Let $(s_{k})$, $k=0,1,2,\ldots$, be a sequence of complex numbers. Let $(w_{k})$ be the sequence given by
$w_{k}=\frac{1}{k+1}\sum_{j=0}^{k-1}s_{j},\quad k=0,1,2,\ldots$ (8.26)
If $(s_{k})$ is a self-inverse sequence, then $(w_{k})$ is an anti-self-inverse sequence while if $(s_{k})$ is an anti-self-inverse sequence, then $(w_{k})$ is a self-inverse sequence.
The implication of Corollary 8.11 is that it is always possible to construct an anti-self-inverse sequence from a given self-inverse sequence and vice versa.
Example 8.4.
By choosing $s_{k}=1/(k+1)$ in (8.26), we deduce that the sequence $(w_{k})$, where
$w_{k}=\frac{H_{k}}{k+1},\quad k=0,1,2,\ldots,$ (8.27)
is an anti-self-inverse sequence. Similarly, $s_{k}=H_{k-1+\delta_{k0}}$ yields $(w_{k})$, where
$w_{k}=\frac{k-1}{k+1}(H_{k-1}-1),\quad k=1,2,3,\ldots,$ (8.28)
as an anti-self-inverse sequence. The choice $s_{k}=L_{k}$ in (8.26) establishes the sequence $(w_{k})$ given by
$w_{k}=\frac{L_{k+1}-1}{k+1},\quad k=0,1,2,\ldots,$ (8.29)
as an anti-self-inverse sequence. Taking $s_{k}=F_{k}$ in (8.26) produces the sequence $(w_{k})$ given by
$w_{k}=\frac{F_{k+1}-1}{k+1},\quad k=0,1,2,\ldots,$ (8.30)
as a self-inverse sequence.
Remark 8.3.
Since the sequence $(a_{k})$, $k=0,1,2,\ldots$, where $a_{k}=1/(k+1)$ is a self-inverse sequence, $w_{k}$ given in (8.30) can simply be replaced with
$w_{k}=\frac{F_{k+1}}{k+1},\quad k=0,1,2,\ldots.$ (8.31)
The next result is an immediate consequence of identity (7.48).
Theorem 8.12.
If $(s_{k})$, $k=0,1,2,\ldots$, is a self-inverse sequence, then the sequence $(w_{k})$, where
$w_{k}=2^{-k}\sum_{j=0}^{k-1}\binom{k}{j}s_{j},$ (8.32)
is an anti-self-inverse sequence, and vice versa.
Theorem 8.13.
Let $j$ be an arbitrary non-negative integer. For $k=0,1,2,\ldots$, we have a self-inverse sequence $(a_{k})$, where
$a_{k}=2^{-k}\binom{k}{j}\left(1+(-1)^{j}\right),$ (8.33)
and an anti-self-inverse sequence $(b_{k})$ given by
$b_{k}==2^{-k}\binom{k}{j}\left(1-(-1)^{j}\right).$ (8.34)
Proof.
On account of (3.46) on page 3.46, use
$s_{k}=\binom{k}{j}\text{ and }\sigma_{k}=(-1)^{j}\delta_{kj},$
in (7.47) to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{k}}{j}}=(-1)^{j}2^{-n}\binom{{n}}{j}$, (8.35)
from which (8.33) and (8.34) follow, in view of Theorem 8.1. ∎
Remark 8.4.
Since $j$ is an arbitrary integer, the self-inverse sequence $(a_{k})$ and the anti-self-inverse sequence $(b_{k})$ in Theorem 8.13 are given more briefly by
$a_{k}=2^{-k}\binom{k}{2j},\quad b_{k}=2^{-k}\binom{k}{2j-1}$. (8.36)
Theorem 7.23 yields the result stated in the next theorem.
Theorem 8.14.
Let $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, be a binomial transform pair of the first kind. Then, for $k=1,2,3,\ldots$, we have a self-inverse sequence $(a_{k})$ given by
$a_{k}=k\left(s_{k}+\sigma_{k}-\sigma_{k-1}\right),$ (8.37)
and an anti-self-inverse sequence $(b_{k})$ given by
$b_{k}=k\left(s_{k}-\sigma_{k}+\sigma_{k-1}\right).$ (8.38)
Corollary 8.15.
If $(s_{k})$, $k=1,2,3,\ldots,$ is a self-inverse sequence, then the sequence $(b_{k})$ given by
$b_{k}=ks_{k-1}$ (8.39)
is an anti-self-inverse sequence. Similarly, if $(s_{k})$, $k=1,2,3,\ldots$, is an anti-self-inverse sequence, then the sequence $(a_{k})$ given by
$a_{k}=ks_{k-1}.$ (8.40)
is a self-inverse sequence.
Example 8.5.
Choosing $s_{k}=L_{k}=\sigma_{k}$, for $k=0,1,2,\ldots$, in (8.37) and (8.38) yields the self-inverse sequence $(a_{k})$, where
$a_{k}=kF_{k-1},$ (8.41)
and an anti-self-inverse sequence, $(b_{k})$, given by
$b_{k}=kL_{k-1}.$ (8.42)
The choice $s_{k}=1/(k+1)=\sigma_{k}$ in (8.37) produces the self-inverse sequence $(a_{k})$, where
$a_{k}=\frac{k-1}{k+1},\quad k=1,2,3,\ldots$ (8.43)
With $s_{k}=H_{k-1}$, $k=1,2,3,\ldots$, in (8.37) we obtain the anti-self-inverse sequence $(b_{k})$, where
$b_{k}=kH_{k-2},\quad k=2,3,4,\ldots$ (8.44)
Theorem 8.16.
Let $m$, $n$ and $r$ be non-negative integers. Let $(w_{k})$, $k=m,m+1,m+2,\ldots$, be a sequence of complex numbers. If $(w_{k})$ is a self-inverse sequence, then
$\sum_{k=r+m}^{n}{(-1)^{k}\binom{{n}}{k}\binom{k}{r}w_{k-r}}=(-1)^{r}\binom{n}{r}w_{n-r},$ (8.45)
while if $(w_{k})$ is an anti-self-inverse sequence, then
$\sum_{k=r+m}^{n}{(-1)^{k}\binom{{n}}{k}\binom{k}{r}w_{k-r}}=(-1)^{r-1}\binom{n}{r}w_{n-r}.$ (8.46)
Thus, if $(w_{k})$ is a self-inverse sequence, then the sequences $(t_{k})$ and $(\tau_{k})$, $k=r+m,r+m+1,r+m+2,\ldots$, where
$t_{k}=\binom{k}{r}w_{k-r}\text{ and }\tau_{k}=(-1)^{r}\binom{k}{r}w_{k-r},$ (8.47)
are a binomial-transform pair of the first kind. Similarly, if $(w_{k})$ is an anti-self-inverse sequence, then the sequences $(t_{k})$ and $(\tau_{k})$, where
$t_{k}=\binom{k}{r}w_{k-r}\text{ and }\tau_{k}=(-1)^{r-1}\binom{k}{r}w_{k-r},$ (8.48)
are a binomial-transform pair of the first kind.
Proof.
Repeated application of (8.39) and (8.40) to the sequence $(w_{k})$. ∎
Example 8.6.
From the Fibonacci sequence $(F_{k})$, $k=0,1,2,\ldots$, and (8.48), we find that the sequences $(t_{k})$ and $(\tau_{k})$, $k=r,r+1,r+2,\ldots$, where
$t_{k}=\binom{k}{r}F_{k-r}\text{ and }(-1)^{r-1}\binom{k}{r}F_{k-r},$ (8.49)
are a binomial-transform pair of the first kind.
The harmonic number self-inverse sequence $H_{k-1}$, $k=1,2,3,\ldots$, in conjunction with (8.47) produces the sequences $(t_{k})$ and $(\tau_{k})$, $k=r+1,r+2,r+3,\ldots$, where
$t_{k}=\binom{k}{r}H_{k-r-1}\text{ and }\tau_{k}=(-1)^{r}\binom{k}{r}H_{k-r-1},$ (8.50)
as a binomial-transform pair of the first kind.
From the self-inverse sequence $(a_{k})$ given in (8.36) and the result stated in (8.47), we find that the sequences $(t_{k})$ and $(\tau_{k})$, $k=r,r+1,r+2,\ldots$, given by
$t_{k}=\binom{{k}}{r}2^{-k}\binom{{k-r}}{{2j}}\text{ and }\tau_{k}=(-1)^{r}\binom{{k}}{r}2^{-k}\binom{{k-r}}{{2j}}$, (8.51)
constitute a binomial-transform pair of the first kind.
9 Various extensions
The results in this section extend some of the binomial convolution identities and other identities derived in Sections 24.
The result stated in the next theorem extends (3.41).
Theorem 9.1.
Let $j$ and $n$ be non-negative integers such that $j\leq n$. Let $\{(s_{k}),(\sigma_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. Then
$\sum_{k=0}^{n}{(-1)^{k}\binom{n-j}{k}2^{n-k}s_{k}}=2^{j}\sum_{k=0}^{n}{\binom{n-j}{k}\sigma_{k}}.$ (9.1)
Proof.
By writing (3.45) in the equivalent form
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{j}2^{k}}=(-1)^{n}\binom{{n}}{j}2^{j},$
we see that $\binom{k}{j}2^{k}$ and $(-1)^{k}\binom{k}{j}2^{j}$ are a binomial pair of the first kind. Using
$t_{k}=\binom{k}{j}2^{k}\text{ and }\tau_{k}=(-1)^{k}\binom{k}{j}2^{j},$ (9.2)
in (2.1) gives (9.1). ∎
Theorem 9.2.
Let $j$ and $n$ be non-negative integers. Let $\{(s_{k}),(\sigma_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. Then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{n-k}}{j}2^{k}s_{k}}=\sum_{k=0}^{n}{(-1)^{j-k}\binom{{n}}{k}\binom{{n-k}}{j}2^{k}\sigma_{k}}.$ (9.3)
Proof.
From (8.35), identify the binomial-transform pair of the first kind:
$t_{k}=2^{-k}\binom{{k}}{j}\text{ and }\tau_{k}=(-1)^{j}2^{-k}\binom{{k}}{j}$,
and plug these into (2.1) to obtain (9.3). ∎
Corollary 9.3.
Let $j$ and $n$ be non-negative integers having different parity. If $(s_{k})$ is a self-inverse sequence, that is $\sigma_{k}=s_{k}$, $k=0,1,2,\ldots$, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{n-k}}{j}2^{k}s_{k}}=0.$ (9.4)
In particular, if $n$ is an odd integer and $(s_{k})$ is a self-inverse sequence, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}2^{k}s_{k}}=0.$ (9.5)
Corollary 9.4.
Let $j$ and $n$ be non-negative integers having the same parity. If $(s_{k})$ is an anti-self-inverse sequence, that is $\sigma_{k}=-s_{k}$, $k=0,1,2,\ldots$, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{n-k}}{j}2^{k}s_{k}}=0.$ (9.6)
In particular, if $n$ is an even integer and $(s_{k})$ is an anti-self-inverse sequence, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}2^{k}s_{k}}=0.$ (9.7)
Theorem 9.5.
Let $n$ be a non-negative integer. If $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ and $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the second kind, then
$\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}\bar{s}_{k}\bar{\tau}_{n-k}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}\bar{t}_{n-k}\sum_{j=0}^{k}{\binom{{k}}{j}\bar{\sigma}_{j}}}.$ (9.8)
Proof.
From (3.41), we can identity the following binomial-transform pair of the second kind:
$\bar{a}_{k}=2^{-k}\bar{s}_{k}\text{ and }\bar{\alpha}_{k}=2^{-k}\sum_{j=0}^{k}{\binom{{k}}{j}\bar{\sigma}_{j}}.$
Replace $\bar{s}_{k}$ with $\bar{a}_{k}$ and $\bar{\sigma}_{k}$ with $\bar{\alpha}_{k}$ in (3.25). This completes the proof. ∎
Theorem 9.6.
Let $n$ be a non-negative integer. If $\{(s_{k}),(\sigma_{k})\}$ and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the first kind, then
$\sum_{k=1}^{n}{(-1)^{n-k-1}\binom{{n}}{k}s_{k-1}t_{n-k}}=\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\tau_{n-k}\sum_{j=1}^{k}{\sigma_{j-1}}}.$ (9.9)
Proof.
In view of the result stated in Theorem 7.6, let
$a_{k}=(1-\delta_{k0})s_{k-1+\delta_{k0}}\text{ and }\alpha_{k}=-\sum_{j=1}^{k}{\sigma_{j-1}},\quad k=0,1,2,\ldots.$
Now, replace $s_{k}$ with $a_{k}$ and $\sigma_{k}$ with $\alpha_{k}$ in (2.1). ∎
We bring this section to a close by giving another generalization of (2.1) based on the result from Corollary 7.24.
Theorem 9.7.
Let $n$ be a non-negative integer. Let $\{(s_{k}),(\sigma_{k})\}$ and $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, are binomial-transform pairs of the first kind. Let $S_{0}(k)=\sigma_{k}$ and $S_{m}(k)=k(S_{m-1}(k)-S_{m-1}(k-1))$ for every positive integer $m$. Then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}k^{m}s_{k}t_{n-k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}S_{m}(k)\tau_{n-k}}.$ (9.10)
In particular,
$\sum_{k=1}^{n}{(-1)^{n-k}\binom{{n}}{k}ks_{k}t_{n-k}}=\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}k(\sigma_{k}-\sigma_{k-1})\tau_{n-k}},$ (9.11)
with the special value
$\sum_{k=1}^{n}{(-1)^{n-k-1}\binom{n}{k}t_{n-k}}=\sum_{k=1}^{n}{(-1)^{k}\binom{n}{k}\tau_{n-k}},$ (9.12)
obtained with $s_{k}=-1/k$ and $\sigma_{k}=H_{k}$.
Proof.
Replace $s_{k}$ with $k^{m}s_{k}$ and $\sigma_{k}$ with $S_{m}(k)$ in (2.1). ∎
10 Polynomial identities involving a binomial transform pair
We have already encountered the identities stated in the next theorem; they are variations on (2.3) and (3.3).
Theorem 10.1.
If $n$ is a non-negative integer and $y$ is a complex variable, then
$\sum_{k=0}^{n}{\binom{{n}}{k}s_{n-k}y^{k}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\sigma_{n-k}\left({1+y}\right)^{k}},$ (10.1)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{s}_{n-k}y^{k}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{\sigma}_{n-k}\left({1+y}\right)^{k}}.$ (10.2)
The implication of Theorem 10.1 is that every binomial-transform pair has associated with it a polynomial identity.
Remark 10.1.
Identity (10.2) is equivalent to Boyadzhiev’s [7, Corollary 1].
11 Binomial transform identities associated with polynomial identities of a certain type
Polynomial identities of the following form abound in the literature:
$\sum_{k=0}^{n}{f(k)t^{p(k)}}=\sum_{k=0}^{m}{g(k)\left({1-t}\right)^{q(k)}};$ (11.1)
where $m$ and $n$ are non-negative integers, $t$ is a complex variable, $p(k)$ and $q(k)$ are sequences of integers, and $f(k)$ and $g(k)$ are sequences of complex numbers.
The binomial identities stated in Theorems 11.1 and 11.2 are readily derived using the operators $\mathcal{L}_{y}$ and $\mathcal{M}_{y}$.
Theorem 11.1.
Let a polynomial identity have the form stated in (11.1). If $\{(s_{k}),(\sigma_{k})\}$ is a binomial-transform pair of the first kind, then
$\sum_{k=0}^{n}{f(k)s_{p(k)}}=\sum_{k=0}^{m}{g(k)\sigma_{q(k)}}.$ (11.2)
Theorem 11.2.
Let a polynomial identity have the form stated in (11.1). If $\{(\bar{s}_{k}),(\bar{\sigma}_{k})\}$ is a binomial-transform pair of the second kind, then
$\sum_{k=0}^{n}{(-1)^{p(k)}f(k)\bar{s}_{p(k)}}=\sum_{k=0}^{m}{g(k)\bar{\sigma}_{q(k)}}.$ (11.3)
Lemma 11.1.
Sun [27, Lemma 3.1] If $m$, $n$ and $r$ are non-negative integers and $t$ is a complex variable, then
$\sum_{k=0}^{n}{(-1)^{k-r}\binom{{n}}{k}\binom{{k+m}}{r}t^{k+m-r}}=\sum_{k=0}^{m}{(-1)^{k}\binom{{m}}{k}\binom{{k+n}}{r}\left({1-t}\right)^{n+k-r}}.$ (11.4)
Example 11.1.
In (11.4) we can identify
$f(k)=(-1)^{k-r}\binom{{n}}{k}\binom{{k+m}}{r},\quad g(k)=\binom{{m}}{k}\binom{{k+n}}{r}$, (11.5)
and
$p(k)=k+m-r\mbox{ and }q(k)=n+k-r.$ (11.6)
Using these in Theorem 11.1 gives
$\sum_{k=0}^{n}{(-1)^{k-r}\binom{{n}}{k}\binom{{k+m}}{r}s_{k+m-r}}=\sum_{k=0}^{m}{(-1)^{k}\binom{{m}}{k}\binom{{k+n}}{r}\sigma_{n+k-r}},$ (11.7)
for every binomial-transform pair $\{(s_{k}),(\sigma_{k})\}$ of the first kind.
Remark 11.1.
Identity (11.7) is Chen’s Theorem 3.2 [9].
12 Conversion between binomial transform identities
If a binomial transform identity involves a binomial-transform pair of the first kind, it is straightforward to convert it to an identity involving a binomial-transform pair of the second kind; and vice versa. An identity involving $\{s(k),\sigma(k)\}$, $k=0,1,2,\dots$, can be converted to that involving $\{\bar{s}(k),\bar{\sigma}(k)\}$ by choosing $s(k)=x^{k}$ and $\sigma(k)=(1-x)^{k}$, replacing $x$ with x-x and then operating on the resulting identity with $\mathcal{M}_{x}$. Similarly, an identity involving $\{\bar{s}(k),\bar{\sigma}(k)\}$ can be converted to that involving $\{s(k),\sigma(k)\}$ by choosing $\bar{s}(k)=x^{k}$ and $\bar{\sigma}(k)=(1+x)^{k}$, replacing $x$ with x-x and then operating on the resulting identity with $\mathcal{L}_{x}$.
Example 12.1.
Let $\{(\bar{t}_{k}),(\bar{\tau}_{k})\}$, $k=0,1,2,\ldots$, be a binomial-transform pair of the second kind. Chen’s main result [9, Theorem 3.1] (after Gould and Quaintance’s simplification [13, Equation (10)]) for non-negative integers $m$, $n$ and $s$ is
$\sum_{k=0}^{m}{\binom{{m}}{k}\binom{{n+k+s}}{s}^{-1}\bar{t}_{n+k+s}}$
$\qquad=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{m+k+s}}{s}^{-1}\bar{\tau}_{m+k+s}}$
$\qquad\qquad+\sum_{k=0}^{s-1}{\frac{{(-1)^{n+s-k}s}}{{m+n+s-k}}\binom{{s-1}}{k}\binom{{m+n+s-k-1}}{n}^{-1}\bar{\tau}_{k}}.$ (12.1)
In order to convert (12.1) to an identity for the binomial-transform pair of the first kind, $\{(t_{k}),(\tau_{k})\}$, $k=0,1,2,\ldots$, we choose $\bar{t}_{k}=x^{k}$ and $\bar{\tau}_{k}=(1+x)^{k}$ and replace $x$ with x-x to obtain
$\sum_{k=0}^{m}{(-1)^{k}\binom{{m}}{k}\binom{{n+k+s}}{s}^{-1}x^{n+k+s}}$
$\qquad=(-1)^{s}\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{m+k+s}}{s}^{-1}(1-x)^{m+k+s}}$
$\qquad\qquad+\sum_{k=0}^{s-1}{\frac{{(-1)^{k}s}}{{m+n+s-k}}\binom{{s-1}}{k}\binom{{m+n+s-k-1}}{n}^{-1}(1-x)^{k}}.$
Operating on the above equation with the linear operator $\mathcal{L}_{x}$ now gives
$\sum_{k=0}^{m}{(-1)^{k}\binom{{m}}{k}\binom{{n+k+s}}{s}^{-1}t_{n+k+s}}$
$\qquad=(-1)^{s}\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{m+k+s}}{s}^{-1}\tau_{m+k+s}}$
$\qquad\qquad+\sum_{k=0}^{s-1}{\frac{{(-1)^{k}s}}{{m+n+s-k}}\binom{{s-1}}{k}\binom{{m+n+s-k-1}}{n}^{-1}\tau_{k}}.$ (12.2)
Example 12.2.
Another example, Chen [9, Theorem 3.2] is
$\sum_{k=0}^{m}{\binom{{m}}{k}\binom{{n+k}}{s}\bar{t}_{n+k-s}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\binom{{m+k}}{s}\bar{\tau}_{m+k-s}};$ (12.3)
whose corresponding version for a binomial-transform pair of the first kind is
$\sum_{k=0}^{m}{(-1)^{s-k}\binom{{m}}{k}\binom{{n+k}}{s}t_{n+k-s}}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{m+k}}{s}\tau_{m+k-s}}.$ (12.4)
Example 12.3.
Gould and Quaintance [13, Theorem 3] gave the following identity involving a binomial transform pair of the second kind:
$\sum_{k=0}^{s}{\binom{{s}}{k}\binom{{m+n+s-k}}{m}^{-1}\frac{{\bar{t}_{k}}}{{m+n+s+1-k}}}$
$\qquad=\sum_{k=0}^{s}{\binom{{s}}{k}\binom{{m+n+s-k}}{n}^{-1}\frac{{(-1)^{s-k}\bar{\tau}_{k}}}{{m+n+s+1-k}}},$ (12.5)
which is valid for every non-negative integer $s$ and all complex numbers $m$ and $n$ excluding the set of negative integers.
The corresponding result for a binomial-transform pair of the first kind is
$\sum_{k=0}^{s}{(-1)^{k}\binom{{s}}{k}\binom{{m+n+s-k}}{m}^{-1}\frac{{t_{k}}}{{m+n+s+1-k}}}$
$\qquad=\sum_{k=0}^{s}{\binom{{s}}{k}\binom{{m+n+s-k}}{n}^{-1}\frac{{(-1)^{s-k}\tau_{k}}}{{m+n+s+1-k}}}.$ (12.6)
Remark 12.1.
In view of the foregoing, an identity involving a binomial-transform pair of the first kind $\{s_{k},\sigma_{k}\}$, $k=0,1,2,\ldots$, can always be converted to a binomial-transform pair identity of the second kind $\{\bar{s}_{k},\bar{\sigma}_{k}\}$ by doing
$s_{k}\to(-1)^{k}\bar{s}_{k},\quad\sigma_{k}\to\bar{\sigma}_{k}.$
Similarly, conversion of a second kind transform pair identity to a first kind transform pair is achieved through
$\bar{s}_{k}\to(-1)^{k}s_{k},\quad\bar{\sigma}_{k}\to\sigma_{k}.$
13 A table of binomial-transform pairs
Many identities involving binomial transforms are known in the literature; ready examples can be found in the books by Gould [14] and Boyadzhiev [7] and in the numerous references therein. In this section we add to the growing list of binomial-transform pairs.
Lemma 13.1.
If $r$ is a non-negative integer, then
$\int_{0}^{1}{\left({1+x}\right)^{r}\ln\left({1+x}\right)dx}=\frac{{1-2^{r+1}}}{{\left({r+1}\right)^{2}}}+\frac{{2^{r+1}}}{{r+1}}\ln 2$ (13.1)
and
[…]
Proof.
Identity (13.1) is obvious. A simple change of variable in the known integral
$\int_{0}^{1}{x^{r}\ln\left({1-x}\right)dx}=-\frac{{H_{r+1}}}{{r+1}}$ (13.3)
gives
$\int_{0}^{1}{x^{r}\ln\left({1+x}\right)dx}=\frac{{H_{r+1}-H_{(r+1)/2}}}{{r+1}},\quad r\in\mathbb{C}\setminus\mathbb{Z}^{-},$ (13.4)
of which (13.2) is a special case, since
$H_{k+1/2}=2\,O_{k+1}-2\ln 2.$
Lemma 13.2.
Let $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, be a binomial-transform pair of the first kind. Let $\mathcal{N}_{k}$ be the linear operator defined by
$\mathcal{N}_{k}(a_{k})=\sum_{j=0}^{k}{(-1)^{j}\binom{{k}}{j}a_{j}}=\alpha_{k}.$
Then
$\mathcal{N}_{k}(\alpha_{k})=a_{k},$ (13.5)
$\mathcal{N}_{k}^{2}(\alpha_{k})=\alpha_{k},$ (13.6)
$\mathcal{N}_{k}^{2}(a_{k})=a_{k},$ (13.7)
with the inverse relations
$\mathcal{N}_{k}^{-1}(a_{k})=\alpha_{k}$ (13.8)
and
$\mathcal{N}_{k}^{-1}(\alpha_{k})=a_{k}.$ (13.9)
Proposition 13.1.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{k}}}{{\left({k+1}\right)^{2}}}}=\frac{{O_{\left\lceil{\left({n+1}\right)/2}\right\rceil}}}{{n+1}}$ (13.10)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{k}}}{{\left({k+1}\right)^{3}}}}$ $=\frac{1}{4}\frac{{H_{n+1}^{2}+H_{n+1}^{(2)}}}{{n+1}}-\frac{1}{8}\frac{{H_{\left\lceil{n/2}\right\rceil}^{2}+H_{\left\lceil{n/2}\right\rceil}^{(2)}}}{{n+1}}+\frac{1}{2}\frac{{O_{\left\lceil{(n+1)/2}\right\rceil}^{2}+O_{\left\lceil{(n+1)/2}\right\rceil}^{(2)}}}{{n+1}}$
$\qquad-\frac{1}{2}\frac{1}{{n+1}}\sum_{k=0}^{n}{(-1)^{k}\frac{{H_{k+1}}}{{k+1}}}.$ (13.11)
Proof.
Write (7.32) as
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n+1}}{{k+1}}\left({1+x}\right)^{k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{x^{2k}}-\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{x^{2k-1}},$
and termwise integrate from 0 to 11, using Lemma 13.1, to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n+1}}{{k+1}}\frac{{1-2^{k+1}}}{{\left({k+1}\right)^{2}}}}+\ln 2\sum_{k=0}^{n}{(-1)^{k}\binom{{n+1}}{{k+1}}\frac{{2^{k+1}}}{{k+1}}}$
$\qquad=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\frac{{H_{2k+1}-2O_{k+1}+2\ln 2}}{{2k+1}}}-\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{\frac{{H_{2k}-H_{k}}}{{2k}}}$
Comparing rational and irrational parts leads to
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n+1}}{{k+1}}\frac{{2^{k}}}{{k+1}}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\frac{1}{{2k+1}}}=O_{\left\lfloor{(n+2)/2}\right\rfloor}$ (13.12)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n+1}}{{k+1}}\frac{{1-2^{k+1}}}{{\left({k+1}\right)^{2}}}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\frac{{H_{2k+1}-2O_{k+1}}}{{2k+1}}}-\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{\frac{{H_{2k}-H_{k}}}{{2k}}}.$ (13.13)
Identity (13.12) immediately gives (13.10) since
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{1}{{\left({k+1}\right)^{3}}}}=\frac{1}{2}\frac{{H_{n+1}^{2}+H_{n+1}^{(2)}}}{{n+1}},$ (13.14)
$\sum_{k=1}^{n}{\frac{{H_{k}}}{k}}=\frac{1}{2}\left({H_{n}^{2}+H_{n}^{(2)}}\right),$ (13.15)
and
$\sum\limits_{k=1}^{n}{\frac{{O_{k}}}{{2k-1}}}=\frac{1}{2}\left({O_{n}^{2}+O_{n}^{(2)}}\right),$ (13.16)
identity (13.13) reduces to
$\frac{1}{2}\left({H_{n+1}^{2}+H_{n+1}^{(2)}}\right)-2(n+1)\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{k}}}{{\left({k+1}\right)^{3}}}}$
$\qquad=\sum\limits_{k=0}^{n}{(-1)^{k}\frac{{H_{k+1}}}{{k+1}}}+\frac{1}{4}\left({H_{\left\lceil{n/2}\right\rceil}^{2}+H_{\left\lceil{n/2}\right\rceil}^{(2)}}\right)-O_{\left\lceil{(n+1)/2}\right\rceil}^{2}-O_{\left\lceil{(n+1)/2}\right\rceil}^{(2)},$
and hence (13.1). ∎
Proposition 13.2.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{(-1)^{k-1}\binom{{n}}{k}\frac{{2^{k}}}{{k^{2}}}}$ $=\frac{1}{2}\left({H_{n}^{2}+H_{n}^{(2)}}\right)-\frac{1}{4}\left({H_{\left\lfloor{n/2}\right\rfloor}^{2}+H_{\left\lfloor{n/2}\right\rfloor}^{(2)}}\right)+O_{\left\lceil{n/2}\right\rceil}^{2}+O_{\left\lceil{n/2}\right\rceil}^{(2)}$
$\qquad-\sum_{k=1}^{n}{(-1)^{k-1}\frac{{H_{k}}}{k}}.$ (13.17)
Proof.
Application of Theorem 7.1 to (13.1). ∎
Proposition 13.3.
If $n$ is a non-negative integer, then
[…]
and
[…]
Proof.
The following identity is known [1]:
[…]
where $v$ is a real number.
Identities (13.18), (13.19) and (13.20) are respective evaluations of (13.21) at $v=0$, $v=1$ and $v=2$. Note that
$\binom{{2k+1}}{k}=\frac{{2k+1}}{{k+1}}\binom{{2k}}{k}$ $=\left({2k+1}\right)C_{k},$ (13.22)
$\binom{{2k+1}}{{\left({2k+1}\right)/2}}$ $=\binom{{2k+1}}{k}^{-1}\frac{{2^{2(2k+1)}}}{{\pi(k+1)}},$ (13.23)
$\binom{{k+1}}{{1/2}}=\binom{{2\left({k+1}\right)}}{{k+1}}^{-1}\frac{{2^{2(k+1)+1}}}{\pi}$ $=\frac{{2^{2(k+1)+1}}}{{2\pi(2k+1)C_{k}}},$ (13.24)
$\binom{{n/2+1/2}}{{1/2}}=\binom{{n/2+1/2}}{{n/2}}=\frac{{n+1}}{{2^{n}}}\binom{{n}}{{n/2}}$, (13.25)
and
$\binom{{2(k+1)}}{{k+1}}=2(2k+1)C_{k}.$ (13.26)
Remark 13.1.
Generalized binomial coefficient identities (13.22)–(13.25) and similar ones can be derived from their definition (1.3) on page 1.3; many of these are listed toward the end of Gould’s book [14].
Remark 13.2.
Identity (13.20) was first discovered by Suleiman and Sury [25].
Proposition 13.4.
If $n$ is a non-negative integer, then
[…]
and
[…]
Proof.
Differentiate (13.21) with respect to $v$ to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{2k+v}}{{(2k+v)/2}}\binom{{k+v}}{{v/2}}^{-1}\left({2\left({H_{2k+v}-H_{k+v}}\right)+H_{v/2}-H_{k+v/2}}\right)}$
$\qquad=\begin{cases}-2^{-n}\binom{{n}}{{n/2}}\binom{{(n+v)/2}}{{v/2}}^{-1}\left({H_{(n+v)/2}-H_{v/2}}\right),&\text{if $n$ is even;}\\ 0,&\text{if $n$ is odd.}\end{cases}$ (13.30)
Setting $v=0$ in (13) gives (13.27). To prove (13.28), set $v=1$ in (13) and use identities (13.22)–(13.25) and the fact that
$H_{1/2}-H_{k+1/2}=2-2\ln 2-(2O_{k+1}-2\ln 2)=2(1-O_{k+1}),$
to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{2^{k}}{k+1}\left(H_{2k+1}-H_{k+1}-O_{k+1}\right)}+\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{k}}}{{k+1}}}$
[…]
which, on account of (13.19) and the fact that
$H_{2k+1}-O_{k+1}-H_{k+1}=\frac{1}{2}H_{k}-H_{k+1}=-\frac{1}{2}H_{k}-\frac{1}{k+1}$
simplifies to
[…]
that is
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{k-1}H_{k}}}{{k+1}}}+\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{k}}}{{\left({k+1}\right)^{2}}}}=\begin{cases}\dfrac{{O_{(n+2)/2}}}{{n+1}},&\text{if $n$ is even;}\\ 0,&\text{if $n$ is odd;}\end{cases}$
from which, together with (13.10), we finally obtain (13.28). Identity (13.29) follows from setting $v=2$ in (13). ∎
Remark 13.3.
Identity (13.27) corresponds to a simplification of [21, Equation (8)]. Identity (13.28) was first derived by Rathie and Campbell [21].
Proposition 13.5.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2k+1}}{{k+1}}\,2^{-k}C_{k}\left({H_{k+1}-O_{k+1}}\right)}$ (13.31)
[…]
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{\left({2k+1}\right)\left({2k+3}\right)}}{{\left({k+1}\right)\left({k+2}\right)^{2}}}\,2^{-k}C_{k}\left({H_{k+2}-O_{k+2}}\right)}$
[…]
Proof.
Application of Theorem 7.3 to (13.27). We used (13.26) and the Catalan number recurrence relation
$C_{n+1}=\frac{{2\left({2n+1}\right)}}{{n+2}}\,C_{n}.$
The next results come from the Knuth-Boyadzhiev identity [6]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}H_{k}x^{k}}=\left({1-x}\right)^{n}H_{n}-\sum_{k=1}^{n}{\frac{{\left({1-x}\right)^{n-k}}}{k}}$ (13.34)
and the following identity due to Batır and Sofo [3]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\bar{H}_{k}x^{k}}=\sum_{k=1}^{n}{\frac{{\left({1-x}\right)^{n-k}}}{k}}-\sum_{k=1}^{n}{\frac{{\left({1-x}\right)^{n-k}\left({1+x}\right)^{k}}}{k}},$ (13.35)
where
$\bar{H}_{n}=\sum_{k=1}^{n}{\frac{(-1)^{k-1}}{k}}$
is the $n^{th}$ alternating harmonic number.
It is known that (see, for example, Frontczak [12]):
$\bar{H}_{n}=2O_{\left\lceil{n/2}\right\rceil}-H_{n},$ (13.36)
and also (Sofo [24]):
$\bar{H}_{n}=H_{n}-H_{\lfloor n/2\rfloor},$ (13.37)
so that
$2\bar{H}_{n}=2O_{\lceil n/2\rceil}-H_{\lfloor n/2\rfloor}.$ (13.38)
From (13.34) and (13.35), using (13.36), (13.37), and (13.38), we find
$2\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}O_{\left\lceil{k/2}\right\rceil}x^{k}}=\left({1-x}\right)^{n}H_{n}-\sum_{k=1}^{n}{\frac{{\left({1-x}\right)^{n-k}\left({1+x}\right)^{k}}}{k}}$ (13.39)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}H_{\left\lfloor{k/2}\right\rfloor}x^{k}}=\left({1-x}\right)^{n}H_{n}-2\sum_{k=1}^{n}{\frac{{\left({1-x}\right)^{n-k}}}{k}}+\sum_{k=1}^{n}{\frac{{\left({1-x}\right)^{n-k}\left({1+x}\right)^{k}}}{k}}.$ (13.40)
Proposition 13.6.
If $n$ is a non-negative integer, then
$2\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}O_{\left\lceil{k/2}\right\rceil}}=-\frac{{2^{n}}}{n},$ (13.41)
$2\sum_{k=0}^{n}{\binom{{n}}{k}O_{\left\lceil{k/2}\right\rceil}}=2^{n}H_{n},$ (13.42)
and
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}2^{k-1}H_{k-1}}=\begin{cases}2\,O_{n/2},&\text{if $n$ is even;}\\ 0,&\text{if $n$ is odd.}\end{cases}$ (13.43)
Proof.
Using $x=1$ and x=1x=-1 in (13.39), in turn, produces (13.41) and (13.42). Identity (13.43) is the result of adding (13.41) and (13.42). ∎
Proposition 13.7.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{O_{\left\lceil{(k+1)/2}\right\rceil}}}{{k+1}}}=\frac{{2^{n}H_{n+1}}}{{n+1}},$ (13.44)
and
[…]
Proof.
Identity (13.44) is obtained from (13.42) by the application of Theorem 7.3. Identity (13.45) follows from (13.43) by the same theorem. ∎
Remark 13.4.
Identity (13.28) can also be derived from (13.43) by the application of Theorem 7.3.
Proposition 13.8.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{k}H_{k+1}}=(-1)^{n}\left({O_{\left\lfloor{(n+1)/2}\right\rfloor}+O_{\left\lceil{(n+1)/2}\right\rceil}}\right)$ (13.46)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}k2^{k}H_{k}}=(-1)^{n}2n\left({O_{\left\lfloor{n/2}\right\rfloor}+O_{\left\lceil{n/2}\right\rceil}}\right).$ (13.47)
Proof.
The binomial inversion of (13.42) gives
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{k}H_{k}}=(-1)^{n}2O_{\left\lceil{n/2}\right\rceil}.$ (13.48)
Addition of (13.19) and (13.48) produces
[…]
from which (13.46) follows. Identity (13.47) is obtained from (13.46) by application of Theorem 7.1. ∎
Proposition 13.9.
If $n$ is a non-negative integer, then
$2\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}kO_{\left\lfloor{k/2}\right\rfloor}}=\frac{{n2^{n-1}}}{{n-1}},\quad n\neq 1,$ (13.49)
$2\sum_{k=0}^{n}{\binom{{n}}{k}kO_{\left\lfloor{k/2}\right\rfloor}}=n2^{n-1}H_{n-1},\quad n\neq 1,$ (13.50)
and
$\sum_{k=2}^{n}{(-1)^{k}\binom{{n}}{k}k2^{k-1}H_{k-2}}=\begin{cases}0,&\text{if $n$ is even;}\\ -4nO_{(n-1)/2},&\text{if $n$ is odd.}\end{cases}$ (13.51)
Proof.
Identities (13.49) and (13.50) are obtained from (13.41) and (13.42) by the application of Theorem 7.1. Subtracting (13.49) from (13.50) gives
$\sum_{k=0}^{n}{\binom{{n}}{k}kO_{\left\lfloor{k/2}\right\rfloor}\left({1-(-1)^{k}}\right)}=n2^{n-1}H_{n-2},\quad n\geq 2,$
from which (13.51) follows by binomial inversion. ∎
Proposition 13.10.
If $m$ and $n$ are non-negative integers, then
$2\sum_{k=m}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{m}O_{\left\lceil{k/2}\right\rceil}}$ $=\frac{{(-1)^{m+1}}}{{m+1}}2^{n-m}\binom{n}{m}\sum_{k=0}^{m}{(-1)^{k}\binom{{k+n}}{n}\binom{{k+n}}{{m+1}}^{-1}},\quad n>m,$ (13.52)
$2\sum_{k=m}^{n}{\binom{{n}}{k}\binom{{k}}{m}O_{\left\lceil{k/2}\right\rceil}}$ $=2^{n-m}\binom{{n}}{m}H_{n}+2^{n-m}\binom{{n}}{m}\sum_{k=1}^{m}{\frac{{(-1)^{k-1}}}{k}\binom{{m}}{k}\binom{{n}}{k}^{-1}},$ (13.53)
and
$\sum_{k=1}^{n}{\frac{{(-1)^{k}}}{k}\binom{{n}}{k}\binom{{m}}{k}^{-1}}=H_{m}-2^{n-m+1}\binom{{m}}{n}^{-1}\sum_{k=1}^{m}{\binom{{m}}{k}\binom{{k}}{n}O_{\left\lceil{k/2}\right\rceil}},\quad m\geq n.$ (13.54)
In particular,
$2\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}kO_{\left\lceil{k/2}\right\rceil}}=-2^{n-1}\frac{{n-2}}{{n-1}}$ (13.55)
and
$2\sum_{k=0}^{n}{\binom{{n}}{k}kO_{\left\lceil{k/2}\right\rceil}}=2^{n-1}\left({nH_{n}+1}\right).$ (13.56)
Proof.
Differentiating (13.39) $m$ times with respect to $x$ gives
$2\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{k}}{m}O_{\left\lceil{k/2}\right\rceil}x^{k-m}}$
$\qquad=\left({1+x}\right)^{n-m}\binom{{n}}{m}H_{n}$
$\qquad\qquad-\sum_{k=1}^{n}{\frac{1}{k}\sum_{j=0}^{m}{(-1)^{j}\binom{{k}}{j}\binom{{n-k}}{{m-j}}\left({1-x}\right)^{k-j}\left({1+x}\right)^{n-k-m+j}}}$
Identities (13.52) and (13.53) are obtained by setting x=1x=-1 and $x=1$, in turn, in the above expression. We used
$\binom{{k-m}}{{n-m}}\binom{{k}}{m}=\binom{{k}}{n}\binom{{n}}{m}$.
Identity (13.54) is a rearrangement of (13.53) with an exchange of $m$ and $n$. ∎
Proposition 13.11.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}H_{\left\lfloor{k/2}\right\rfloor}}$ $=\frac{{2^{n}-2}}{n},$ (13.57)
$\sum_{k=0}^{n}{\binom{{n}}{k}H_{\left\lfloor{k/2}\right\rfloor}}$ $=2^{n}H_{n}-2^{n+1}\sum_{k=1}^{n}{\frac{1}{{k2^{k}}}}$ (13.58)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{k}H_{\left\lfloor{k/2}\right\rfloor}}=(-1)^{n}3H_{n}-(-1)^{n}2H_{\left\lfloor{n/2}\right\rfloor}+\sum_{k=1}^{n}{(-1)^{n-k}\frac{{3^{k}}}{k}}.$ (13.59)
Proof.
Set $x=1$, x=1x=-1 and $x=2$, in turn, in (13.40). ∎
Proposition 13.12.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{H_{\left\lceil{k/2}\right\rceil}}}{{k+1}}}=\frac{{2-2^{n+1}}}{{\left({n+1}\right)^{2}}}$ (13.60)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{H_{\left\lceil{(k+1)/2}\right\rceil}}}{{\left({k+1}\right)\left({k+2}\right)}}}=-\frac{{2-2^{n+2}}}{{\left({n+1}\right)\left({n+2}\right)^{2}}}.$ (13.61)
Proof.
Application of Theorem 7.3 to (13.57). ∎
Proposition 13.13.
If $n$ is a non-negative integer, then
$H_{n}=\frac{1}{2}H_{\left\lfloor{n/2}\right\rfloor}+O_{\left\lceil{n/2}\right\rceil}.$ (13.62)
Proof.
Act on (13.57) with $\mathcal{N}_{n}$; making use of (8.5) and (13.41). ∎
Proposition 13.14.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{C_{k-1}}}{{2^{2k}}}}$ $=\frac{1}{2}-\frac{{2n+1}}{{2^{2n+1}}}\binom{{2n}}{n}$, (13.63)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{-2k}C_{k}}}{{k+1}}}$ $=\left({2n+1}\right)\left({2n+3}\right)\frac{{2^{-2n}C_{n}}}{{n+1}}-\frac{2}{{n+1}},$ (13.64)
$\sum_{k=1}^{n}{\binom{{n}}{k}\frac{{C_{k-1}}}{{2^{2k}}}}$ $=\frac{1}{2}-\frac{1}{{2^{2n+1}}}\binom{{2n}}{n}+2^{-2n}\binom{{2n}}{n}\sum_{k=1}^{\left\lfloor{n/2}\right\rfloor}{2^{2k}C_{k-1}\binom{{n}}{{2k}}\binom{{4k}}{{2k}}^{-1}},$ (13.65)
and
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{C_{k}}}{{2^{2k}(k+1)}}}$ $=\frac{{2\left({2n+1}\right)}}{{2^{2n}\left({n+1}\right)}}C_{n}\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{2^{2k}C_{k-1}\binom{{n+1}}{{2k}}\binom{{4k}}{{2k}}^{-1}}$
$\qquad+\frac{2}{{n+1}}-\frac{{2n+1}}{{2^{2n}\left({n+1}\right)}}C_{n}.$ (13.66)
Proof.
Write (3.42) as
$\sum_{k=1}^{n}{\binom{{n}}{k}\binom{{x}}{{k+z}}}=\binom{{n+x}}{{n+z}}-\binom{x}{z}$
and set $z=0$ and $x=1/2$ to obtain
$\sum_{k=1}^{n}{\binom{{n}}{k}\binom{{1/2}}{k}}=\binom{{n+1/2}}{n}-1$
and hence (13.63) since
$\binom{{1/2}}{s}=\frac{{(-1)^{s-1}}}{{2^{2s-1}}}C_{s-1}$
and
$\binom{{s+1/2}}{s}=\frac{{2s+1}}{{2^{2s}}}\binom{{2s}}{s}$.
Setting $y=n$ and $x=1/2$ in the following identity [20, Equation (7.14), p.35]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{x}}{k}\binom{{y}}{{n-k}}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{(-1)^{k}\binom{x}{k}\binom{{y-x}}{{n-2k}}}$,
where $x$ and $y$ are complex numbers, gives
$\sum_{k=1}^{n}{\binom{{n}}{k}\frac{{C_{k-1}}}{{2^{2k}}}}=\frac{1}{2}-\frac{1}{2}\binom{{n-1/2}}{n}+\sum_{k=1}^{\left\lfloor{n/2}\right\rfloor}{\frac{{C_{k-1}}}{{2^{2k}}}\binom{{n-1/2}}{{n-2k}}}$,
and hence (13.65) since
$\binom{{s-1/2}}{s}=2^{-2s}\binom{{2s}}{s}$
and
$\binom{{r-1/2}}{{r-2s}}=2^{4s-2r}\binom{{2r}}{r}\binom{{r}}{{2s}}\binom{{4s}}{{2s}}^{-1}.$
Identities (13.64) and (13.14) follow from (13.63) and (13.65) by application of Theorem 7.3. ∎
Proposition 13.15.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{C_{k}}}{{2^{2k}}}}$ $=\frac{{2n+1}}{{2^{2n}}}C_{n},$ (13.67)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{kC_{k}}}{{2^{2k}}}}$ $=-\frac{{nC_{n}}}{{2^{2n}}},$ (13.68)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{C_{k+1}}}{{2^{2k}}}}$ $=\frac{{C_{n+1}}}{{2^{2n}}},$ (13.69)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\binom{{2k}}{k}}=2^{-2n}\binom{{2n}}{n}$. (13.70)
Proof.
Set $x=1/2$ and $z=1$ in (3.42) to obtain (13.67) since
$\binom{{1/2}}{{s+1}}=\frac{{(-1)^{s}}}{{2^{2s+1}}}C_{s}$
and
$\binom{{s+1/2}}{{s+1}}=\frac{{2s+1}}{{2^{2s+1}}}C_{s}.$
Identities (13.68) and (13.70) follow from (13.67) by Theorem 8.1. Identity (13.69) is obtained from (13.68) by Theorem 7.3. ∎
Remark 13.5.
Identity (13.69) was also reported in [11].
Lemma 13.3.
If $x$ and $z$ are complex numbers that are not negative integers and $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{{k+z}}H_{x-k-z}}=\binom{{n+x}}{{n+z}}\left({H_{x}-H_{n+x}+H_{x-z}}\right)$ (13.71)
and
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{{k+z}}H_{k+z}}=\binom{{n+x}}{{n+z}}\left({H_{x}-H_{n+x}+H_{n+z}}\right).$ (13.72)
Proof.
Differentiate (3.42) with respect to $x$ to obtain
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{{k+z}}\left({H_{x}-H_{x-k-z}}\right)}=\binom{{n+x}}{{n+z}}\left({H_{n+x}-H_{x-z}}\right)$ (13.73)
and hence (13.71) after using (3.42) again. Differentiating (3.42) with respect to $z$ gives
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{x}}{{k+z}}\left({H_{x-k-z}-H_{k+z}}\right)}=\binom{{n+x}}{{n+z}}\left({H_{x-z}-H_{n+z}}\right),$
which in conjunction with (13.71) gives (13.72). ∎
Proposition 13.16.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\binom{{2k}}{k}O_{k}}=-2^{-2n}\binom{{2n}}{n}O_{n}$ (13.74)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\frac{{\left({2k+1}\right)}}{{k+1}}C_{k}O_{k+1}}=2^{-2n}\frac{{\left({2n+1}\right)}}{{n+1}}C_{n}O_{n+1}.$ (13.75)
Proof.
Set x=1/2x=-1/2 and $z=0$ in (13.73) to obtain (13.74). Identity (13.75) is obtained from (13.74) by applying Theorem 7.3. ∎
Remark 13.6.
Identity (13.74) was also derived in [21].
Proposition 13.17.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{C_{k}O_{k}}}{{2^{2k}}}}=2n\frac{{C_{n}}}{{2^{2n}}}-\left({2n+1}\right)\frac{{C_{n}O_{n}}}{{2^{2n}}},$ (13.76)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{kC_{k}O_{k}}}{{2^{2k}}}}=\frac{{nC_{n}}}{{2^{2n}}}\left({O_{n}-2}\right),$ (13.77)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{C_{k+1}O_{k+1}}}{{2^{2k}\left({k+1}\right)}}}=\left({2n+3}\right)\frac{{C_{n+1}O_{n+1}}}{{2^{2n}\left({n+1}\right)}}-\frac{{2C_{n+1}}}{{2^{2n}}},$ (13.78)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{C_{k+1}O_{k+1}}}{{2^{2k}}}}=\frac{{C_{n+1}}}{{2^{2n}}}\left({2-O_{n+1}}\right).$ (13.79)
Proof.
Set $z=1$ and $x=1/2$ in (13.73) to obtain (13.76) after some straightforward manipulation. Identity (13.77) follows from (13.76) by the application of Theorem 7.23. Identities (13.78) and (13.79) come respectively from (13.76) and (13.77) by Theorem 7.3. ∎
Proposition 13.18.
If $m$ and $n$ are positive integers such that $m\geq n$, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{k}\binom{{2\left({m-k}\right)}}{{m-k}}^{-1}\binom{{2k}}{k}^{-1}}=\frac{{m}}{{m+n}}2^{2n}\binom{{m+n}}{n}\binom{{2m}}{m}^{-1}\binom{{2n}}{n}^{-1}$ (13.80)
and
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{k}\binom{{2\left({m-k}\right)}}{{m-k}}^{-1}\binom{{2k}}{k}^{-1}O_{k}}$
$\qquad=\frac{m}{{m+n}}2^{2n-1}\binom{{m+n}}{n}\binom{{2m}}{m}^{-1}\binom{{2n}}{n}^{-1}\left({H_{m-1}-H_{n+m-1}+2O_{n}}\right).$ (13.81)
Proof.
Set x=m1x=m-1 and z=1/2z=-1/2 in (13.72) and equate the respective rational and irrational parts of the resulting identity. ∎
Proposition 13.19.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{C_{k}H_{k+1}}}{{2^{2k}}}}=\frac{{2n+1}}{{2^{2n}}}C_{n}\left({H_{n+1}-2O_{n+1}+2}\right)$ (13.82)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\left({2k+1}\right)\frac{{C_{k}H_{k+1}}}{{2^{2k}}}}=\frac{1}{{2^{2n}}}C_{n}\left({H_{n+1}-2O_{n}}\right).$ (13.83)
Proof.
Using $x=1/2$ and $z=1$ in (13.72) gives (13.82) while x=1/2x=-1/2 and $z=1$ yields (13.83). ∎
Proposition 13.20.
If $n$ is a non-negative iinteger, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2k}}}{{C_{k}}}}=\frac{3}{{\left({2n-1}\right)\left({2n-3}\right)}}$ (13.84)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2k}}}{{\left({2k+1}\right)C_{k}}}}=-\frac{1}{{\left({2n+1}\right)\left({2n-1}\right)}}.$ (13.85)
Proof.
Consider the following identity [1]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+u+v+1}}{{u+1}}^{-1}}=\frac{{u+1}}{{v+1}}\binom{{n+u+v+1}}{{v+1}}^{-1}.$ (13.86)
Setting $v=1$ and u=5/2u=-5/2 gives (13.84) while $v=1$, u=3/2u=-3/2 yields (13.85) since
$\binom{{k-1/2}}{{k+1}}=-\frac{{C_{k}}}{{2^{2k+1}}}$
and
$\binom{{k+1/2}}{{k+1}}=\frac{{2k+1}}{{2^{2k+1}}}C_{k}.$
Proposition 13.21.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{2k}\binom{{2k}}{k}^{-1}}$ $=-\frac{1}{{2n-1}},$ (13.87)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2k}}}{{2k+1}}\binom{{2k}}{k}^{-1}}$ $=\frac{1}{{2n+1}},$ (13.88)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{1}{{k+1}}}=\frac{1}{{n+1}}.$ (13.89)
Proof.
Set $v=0$ and write u1u-1 for $u$ in (13.86) to get
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+u}}{k}^{-1}}=\frac{u}{{n+u}}.$ (13.90)
Evaluate (13.90) at u=1/2u=-1/21/21/211 to obain (13.87), (13.88) and (13.89). ∎
Lemma 13.4.
If $u$ and $v$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+v}}{u}^{-1}H_{k+v}}$ $=\frac{u}{{v-u+1}}\binom{{v+n}}{{v-u+1}}^{-1}\left({H_{u}-H_{n+u-1}+H_{n+v}}\right)$
$\qquad-\frac{1}{{v-u+1}}\,\binom{{v+n}}{{v-u+1}}^{-1}$ (13.91)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+u+v}}{u}^{-1}}H_{k+v}$ $=\frac{u}{{v+1}}\binom{{n+u+v}}{{v+1}}^{-1}\left({H_{u}+H_{v+1}-H_{n+u-1}}\right)$
$\qquad-\frac{{u+v+1}}{{\left({v+1}\right)^{2}}}\,\binom{{n+u+v}}{{v+1}}^{-1}.$ (13.92)
Proof.
Differentiate (13.86) with respect to $u$ and use (13.86) again to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+u+v+1}}{{u+1}}^{-1}H_{k+u+v+1}}$
$\qquad=\frac{{u+1}}{{v+1}}\binom{{n+u+v+1}}{{v+1}}^{-1}\left({H_{u+1}-H_{n+u}+H_{n+u+v+1}}\right)$
$\qquad\qquad-\frac{1}{{v+1}}\binom{{n+u+v+1}}{{v+1}}^{-1}.$ (13.93)
Now replace $u$ with u1u-1 and $v$ with vuv-u; this gives (13.4). Next differentiate (13.86) to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k+u+v+1}}{{u+1}}^{-1}\left({H_{k+u+v+1}-H_{k+v}}\right)}$
$\qquad=\frac{{u+1}}{{\left({v+1}\right)^{2}}}\binom{{n+u+v+1}}{{v+1}}^{-1}+\frac{{u+1}}{{v+1}}\binom{{n+u+v+1}}{{v+1}}^{-1}\left({H_{n+u+v+1}-H_{v+1}}\right).$ (13.94)
Identity (13.4) now follows from (13) and (13) after replacing $u$ with u1u-1. ∎
Proposition 13.22.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\binom{{2k}}{k}H_{k}}$ $=2^{-2n}\binom{{2n}}{n}\left({H_{n}-2O_{n}}\right),$ (13.95)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\binom{{2k}}{k}\left({H_{k}-O_{k}}\right)}$ $=2^{-2n}\binom{{2n}}{n}\left({H_{n}-O_{n}}\right),$ (13.96)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\frac{{2k+1}}{{k+1}}C_{k}\left({H_{k+1}-O_{k+1}}\right)}=-2^{-2n}\frac{{2n+1}}{{n+1}}C_{n}\left({H_{n+1}-O_{n+1}}\right).$ (13.97)
Proof.
Set $v=0$, $u=1/2$ in (13.4) to obtain (13.95). Identity (13.96) follows from (13.95) by applying Theorem 8.1. Identity (13.97) is obtained from (13.96) through the application of Theorem 7.3. ∎
Proposition 13.23.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{2k}\frac{{O_{k}}}{{C_{k}}}}=\frac{{4n\left({4n-5}\right)}}{{\left({2n-1}\right)^{2}\left({2n-3}\right)^{2}}}$ (13.98)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2k}}}{{k+1}}\frac{{O_{k+1}}}{{C_{k+1}}}}=-\frac{{\left({4n-1}\right)}}{{\left({2n+1}\right)^{2}\left({2n-1}\right)^{2}}}.$ (13.99)
Proof.
Set v=1/2v=-1/2, u=3/2u=-3/2 in (13.4) to obtain (13.98), from which identity (13.99) then follows by Theorem 7.3. ∎
Proposition 13.24.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{2k}\binom{{2k}}{k}^{-1}H_{k}}=\frac{{2\left({O_{n-1}-1}\right)}}{{2n-1}}$ (13.100)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{2k}\binom{{2k}}{k}^{-1}\frac{{H_{k+1}}}{{2k+1}}}=-\frac{{\left({O_{n}-1}\right)}}{{\left({n+1}\right)\left({2n+1}\right)}}.$ (13.101)
Proof.
Put $v=0$, u=1/2u=-1/2 in (13.4); this gives (13.100) from which (13.101) follows by Theorem 7.3. ∎
Proposition 13.25.
If $n$ is a non-negative integer and $u$ is a complex number, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{2k}}{k}\binom{{2\left({k+u}\right)}}{{k+u}}^{-1}\binom{{k+u}}{u}^{-1}}=\frac{u\,2^{2n}}{{n+u}}\binom{{2\left({n+u}\right)}}{{n+u}}^{-1}$ (13.102)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{2k}}{k}\binom{{2\left({k+u}\right)}}{{k+u}}^{-1}\binom{{k+u}}{u}^{-1}O_{k}}$
$\qquad=\frac{{u\,2^{2n-1}}}{{n+u}}\binom{{2\left({n+u}\right)}}{{n+u}}^{-1}\left({H_{u}-H_{n+u-1}}\right)-\frac{{2^{2n-1}}}{{n+u}}\binom{{2\left({n+u}\right)}}{{n+u}}^{-1}.$ (13.103)
Proof.
Set v=1/2v=-1/2 in (13.4) and use
$\binom{{k+u-1/2}}{u}=\binom{{2\left({k+u}\right)}}{{k+u}}\binom{{k+u}}{u}2^{-2u}\binom{{2k}}{k}^{-1}.$
Now equate respective rational and irrational parts in the resulting expression. ∎
Proposition 13.26.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{2k}\frac{{H_{k}}}{{C_{k}}}}=\frac{{2n+8-6O_{n}}}{{\left({2n-1}\right)\left({2n-3}\right)}}+\frac{{24\left({n-1}\right)}}{{\left({2n-1}\right)^{2}\left({2n-3}\right)^{2}}}$ (13.104)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{2^{2k}}}{{k+1}}\frac{{H_{k+1}}}{{C_{k+1}}}}=\frac{{3O_{n+1}-n-5}}{{2\left(n+1\right)\left({2n-1}\right)\left({2n+1}\right)}}-\frac{{6n}}{{\left({n+1}\right)\left({2n-1}\right)^{2}\left({2n+1}\right)^{2}}}.$ (13.105)
Proof.
Set $v=1$ and u=3/2u=-3/2 in (13.4) to obtain (13.104) from which (13.105) follows by Theorem 7.3. ∎
Proposition 13.27.
If $n$ is a non-negative integer and $r$ is a complex number that is not a negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{{k+r}}}=\begin{cases}\dfrac{{\sin(\pi r)}}{{\pi\left({r+n}\right)}},&\text{if $r\neq 0$;}\\ 0,&\text{if $r=0$, $n\neq 0$;}\end{cases}$ (13.106)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{{k+r}}H_{k+r}}=\begin{cases}\dfrac{{\sin(\pi r)}}{{\pi\left({r+n}\right)}}\left({H_{r-1}+\dfrac{1}{{r+n}}}\right),&\text{if $r\neq 0$;}\\ -1/n,&\text{if $r=0$, $n\neq 0$.}\end{cases}$ (13.107)
Proof.
Write the identity [14, Entry 1.5]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{p}}{k}}=(-1)^{n}\binom{{p-1}}{n}$ (13.108)
as
$\sum_{k=0}^{n}{(-1)^{k}\binom{{p}}{{n-k}}}=\binom{{p-1}}{n}$
and set $p=r+n$ to obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{r+n}}{{r+k}}}=\binom{{r+n-1}}{n}$, (13.109)
from which identity (13.106) then follows upon using the identity
$\binom{{n+r}}{{k+r}}\binom{{n}}{{n+r}}=\binom{{n}}{k}\binom{{k}}{{k+r}}$,
and the fact that
[…]
Identity (13.107) is obtained by differentiating (13.106) with respect to $r$. ∎
Proposition 13.28.
If $n$ is a positive integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}H_{k}^{2}}$ $=\frac{{H_{n}}}{n}-\frac{2}{{n^{2}}},$ (13.110)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}O_{k}}$ $=-\frac{{2^{2n-1}}}{n}\binom{{2n}}{n}^{-1},$ (13.111)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}O_{k}^{2}}=\frac{{2^{2n-1}}}{n}\binom{{2n}}{n}^{-1}\left({H_{n-1}-O_{n}}\right).$ (13.112)
Proof.
Setting $\ell=0$ in the identity [22]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{n}{k}H_{k+\ell}^{2}}=\frac{1}{n}\binom{n+\ell}{\ell}^{-1}\left({2H_{n-1}-H_{n+l}-H_{l}}\right),\quad\ell\in\mathbb{C}\setminus\mathbb{Z}^{-},$ (13.113)
gives (13.110). Identities (13.111) and (13.112) are obtained by setting $\ell=-1/2$ in (13.113) and equating respective rational and irrational parts in the resulting expression. ∎
Proposition 13.29.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{H_{k+1}^{2}}}{{k+1}}}$ $=-\frac{{H_{n+1}}}{{\left({n+1}\right)^{2}}}+\frac{2}{{\left({n+1}\right)^{3}}},$ (13.114)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{O_{k+1}}}{{k+1}}}$ $=\frac{{2^{2n}}}{{\left({n+1}\right)\left({2n+1}\right)}}\binom{{2n}}{n}^{-1},$ (13.115)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{O_{k+1}^{2}}}{{k+1}}}=-\frac{{2^{2n}}}{{\left({n+1}\right)\left({2n+1}\right)}}\binom{{2n}}{n}^{-1}\left({H_{n}-O_{n+1}}\right).$ (13.116)
Proof.
These identities follow from those in Proposition 13.28 by the application of Theorem 7.3. ∎
Proposition 13.30.
If $n$ is a non-negative integer and $m$ and $p$ are complex numbers, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{p}}{{k+m}}^{-1}}=\frac{{p+1}}{{p+1-n}}\binom{{p-n}}{m}^{-1}$ (13.117)
and
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{p}}{{k+m}}^{-1}H_{k+m}}$ $=\frac{{p+1}}{{p+1-n}}\binom{{p-n}}{m}^{-1}\left({H_{p}-H_{p-n}+H_{m}}\right)$
$\qquad-\frac{n}{{\left({p+1-n}\right)^{2}}}\binom{{p-n}}{m}^{-1}.$ (13.118)
Proof.
Identity (13.117) is the same as (3.28) on page 3.28, with $m$ and $p$ interchanged. From (3.7) and (3.35) we find
$\sum_{k=0}^{n}{\binom{{n}}{k}\binom{{m}}{{p+n-k}}^{-1}H_{m-p-n+k}}$ $=\frac{{m+1}}{{m+1-n}}\binom{{m-n}}{p}^{-1}\left({H_{m}-H_{m-n}+H_{m-n-p}}\right)$
$\qquad-\frac{n}{{\left({m-n+1}\right)^{2}}}\binom{{m-n}}{p}^{-1},$
from which (13.30) follows after first replacing $m$ with $m+p+n$, and then replacing $p$ with pmnp-m-n. ∎
Proposition 13.31.
If $m$ is a non-negative integer and $n$ is a positive integer, then
[…]
and
[…]
Proof.
By (1.6) we identify the binomial-transform pair $(t_{k})$ and $(\tau_{k})$, $k=0,1,2,\ldots$, where
[…]
Now, consider the identity [2]:
$\sum_{j=1}^{k}{j^{m}}=\frac{{B_{m+1}(k+1)-B_{m+1}}}{{m+1}}.$ (13.123)
By Theorem 7.25, we recognize the binomial-transform pair of the first kind $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
$a_{k}=\sum_{j=1}^{k}{j^{m}}=\frac{{B_{m+1}(k+1)-B_{m+1}}}{{m+1}},$ (13.124)
and
[…]
Thus
[…]
and hence (13.119). By Theorem 7.9 we have that
$\frac{{\sum_{j=0}^{k}{t_{j}}}}{{k+1}}\text{ and }\frac{{\tau_{k}}}{{k+1}}$
are a binomial-transform pair of the first kind, with $(t_{k})$ and $(\tau_{k})$ given in (13.122) and hence (13.120). Identity (13.121) follows from (13.120) by the application of Theorem 7.3. ∎
Proposition 13.32.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{kB_{k+1}}{{k+1}}}$ $=\frac{1}{{n+1}}-(-1)^{n}B_{n}+\frac{nB_{n+1}}{{n+1}},$ (13.127)
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{kB_{k+1}}}{{k+1}}}$ $=-\frac{1}{{n+1}}+(-1)^{n}B_{n}+(-1)^{n}\frac{{nB_{n+1}}}{{n+1}},$ (13.128)
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k+2}}}{{k+2}}}$ $=\frac{1}{{\left({n+1}\right)\left({n+2}\right)}}+(-1)^{n}\frac{{B_{n+1}}}{{n+1}}+\frac{{B_{n+2}}}{{n+2}},$ (13.129)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\frac{{B_{k+2}}}{{k+2}}}=\frac{1}{{\left({n+1}\right)\left({n+2}\right)}}+(-1)^{n}\frac{{B_{n+1}}}{{n+1}}+(-1)^{n}\frac{{B_{n+2}}}{{n+2}}.$ (13.130)
Proof.
Apostol [2] derived an identity that can be written in the following equivalent form:
$\sum_{k=2}^{n}{\binom{{n}}{{k-2}}\frac{{B_{k}}}{k}}=\frac{1}{{\left({n+1}\right)\left({n+2}\right)}}-(-1)^{n+1}B_{n+1},$
which produces (13.127) by a shift of the summation index. Identity (13.128) is obtained from (13.127) by acting on both sides with the operator $\mathcal{N}_{n}$ defined in Lemma 13.2 on page 13.2. Identities (13.129) and (13.130) are obtained from (13.127) and (13.128) by Theorem 7.3. ∎
Proposition 13.33.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k+1}}}{{k+1}}}=(-1)^{n+1}\frac{{B_{n+1}}}{{n+1}}-\frac{1}{{n+1}}.$ (13.131)
Proof.
Operating on (13.130) with $\mathcal{N}_{n}$ gives
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k+1}}}{{k+1}}}=-\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k+2}}}{{k+2}}}-\frac{1}{{n+2}}+(-1)^{n}\frac{{B_{n+2}}}{{n+2}},$
from which (13.131) follows after using (13.129). ∎